### v1

# L(M)V-IQL: Multiple Intention Inverse Reinforcement Learning for Animal Behavior Characterization

Abstract
In advancing the understanding of decision-making processes, mathematical models, particularly Inverse Reinforcement Learning (IRL), have proven instrumental in reconstructing animal’s multiple intentions amidst complex behaviors. Given the recent development of a continuous-time multi-intention IRL framework, there has been persistent inquiry into inferring discrete time-varying reward functions with multiple intention IRL approaches. To tackle the challenge, we introduce the Latent (Markov) Variable Inverse Q-learning (L(M)V-IQL) algorithms, a novel IRL framework tailored for accommodating discrete intrinsic rewards. Leveraging an Expectation-Maximization approach, we cluster observed trajectories into distinct intentions and independently solve the IRL problem for each. Demonstrating the efficacy of L(M)V-IQL through simulated experiments and its application to different real mouse behavior datasets, our approach surpasses current benchmarks in animal behavior prediction, producing interpretable reward functions. This advancement holds promise for neuroscience and psychology, contributing to a deeper understanding of animal decision-making and uncovering underlying brain mechanisms.
1 Introduction
The characterization of animal’s decision-making behavior stands as a fundamental objective within the field of behavioral neuroscience (Niv, 2009; Wilson & Collins, 2019). Prior research has formulated a variety of mathematical behavioral models across diverse tasks, encompassing generalized linear models and models based on reinforcement learning (Ashwood et al., 2022b; Beron et al., 2022). Such forward models facilitate the comprehension and comparison of decision-making strategies employed by both human and animal subjects. Additionally, they offer a low-dimensional behavioral representation suitable for regression analysis with neural activities (Hattori et al., 2019; Hamaguchi et al., 2022). Forward models require a empirically defined reward function that guides subjects optimizing their behavior during decision-making. However, defining a comprehensive and suitable reward function can pose challenges in complex behavioral tasks. Alyahyay et al. (2023) introduced a response-preparation task where subjects ought to hold a lever until a cue indicating the release signal. In this task, subjects can receive a binary extrinsic reward from the environment, whereas the intrinsic reward driving behavior such as hunger, thirst, engagement, associated with each timestamp is, however, not obvious to the experimenter. As another example, within a 127-node-labyrinth with a water port at the terminal, Rosenberg et al. (2021) observed that the navigation behavior of water-restricted mice is influenced not solely by the extrinsic water reward but also by intrinsic motivators, including their curiosity to explore the environment.
Inverse reinforcement learning (IRL) (Ng et al., 2000; Arora & Doshi, 2021) is a popular approach to recover a reward function that induces the observed behavior, assuming that the demonstrator was (softly) maximizing its long-term return. Along with the significant successes of IRL in autonomous driving (Kalweit et al., 2020; Nasernejad et al., 2023), robotics (Kumar et al., 2023; Chen et al., 2023), and healthcare domains (Coronato et al., 2020; Chan & van der Schaar, 2021), it appears to be emerging as an valuable tool for constructing mathematical behavior models in neuroscience research, as exemplified by Yamaguchi et al. (2018), Kwon et al. (2020), and Alyahyay et al. (2023). Classic IRL methods seek to identify a single, fixed reward function for a specific scenario. In contrast, Ashwood et al. (2022a) suggested that animal’s goals can evolve over time due to factors like fatigue, satiation, and curiosity. Under this assumption, they proposed the Dynamic Inverse Reinforcement Learning (DIRL) framework, which parametrizes the animal’s reward function as a time-varying linear combination of a small number of spatial reward maps, which are referred to as “goal maps”. By positing the existence of multiple goal maps with time-varying weights, DIRL allows the instantaneous reward function to vary continuously in time. This innovative framework achieved state-of-the-art performance in characterizing animal behavior. Nevertheless, persistent demands have emerged regarding an IRL framework incorporating discrete time-varying reward functions, particularly following the proposal by Ashwood et al. (2022b) that natural behaviors can be represented through a Markov chain characterized by alternating between discrete intentions.
To address this requirement, we propose the novel class of Latent (Markov) Variable Inverse Q-learning (L(M)V-IQL) algorithms, which extend the single reward Inverse Q-learning (IQL) framework from Kalweit et al. (2020) to solve IRL problems accounting for multiple intentions. We formulate an Expectation-Maximization (EM) approach to first cluster animal trajectories into multiple intentions, and then solving the IRL problem independently for each intention, respectively. We theoretically demonstrate that L(M)V-IQL can cover the most common two types of intention transition dynamics: generalized Bernoulli process and Markov process. Finally, we present the application of our framework in 1) a simulated Gridworld environment, 2) real mice navigation trajectories with known environment model from Rosenberg et al. (2021), and 3) real mice decision-making data from a dynamic two-armed bandit task with unknown environment model, showing that our methods outperform state-of-the-art in behavior prediction, demonstrate exceptional aptitude in capturing the intentions of animals, and provide interpretable reward functions corresponding to each identified intention.
2 Related Work
Table 1: Overview of different multiple intention IRL algorithms.
Algorithms Model-free Rewards # Intentions
Time-varying
Rewards
EM-MLIRL (Babes et al., 2011) $\times$ linear known $\times$
DPM-BIRL (Choi & Kim, 2012) $\times$ linear unknown $\times$
MRP/MPO-MC (Dimitrakakis & Rothkopf, 2012) $\times$ linear known discrete
BN-IRL (Michini & How, 2012) $\times$ linear unknown discrete
BNP-IRL (Surana & Srivastava, 2014) $\times$ linear unknown discrete
Meta-AIRL (Gleave & Habryka, 2018) $\checkmark$ non-linear $\times$ $\times$
SEM/MCEM-MIIRL (Bighashdel et al., 2021) $\times$ non-linear unknown $\times$
MI-$\Sigma$-GIRL (Likmeta et al., 2021) $\checkmark$ linear known $\times$
DIRL (Ashwood et al., 2022a) $\times$ non-linear known continuous
L(M)V-IAVI (Ours) $\times$ non-linear known discrete
L(M)V-IQL (Ours) $\checkmark$ non-linear known discrete
Various approaches have been introduced to address multiple intention inverse reinforcement learning problems (Table 1). Notably, several frameworks based on parametric (Babes et al., 2011; Likmeta et al., 2021), or Bayesian non-parametric (Choi & Kim, 2012; Bighashdel et al., 2021) approaches share a similar objective to ours, allowing for multiple agents with distinct reward functions. However, these frameworks do not accommodate single agents with time-varying rewards. Gleave & Habryka (2018) developed a meta adversarial learning method for multi-task IRL problems. While their framework demonstrated high-level performance in real-world applications, it sacrifices theoretical interpretability and heavily relies on exploiting similarities between reward functions across tasks. Built on the Bayesian IRL approach, algorithms from Dimitrakakis & Rothkopf (2012), Michini & How (2012), and Surana & Srivastava (2014) posits that an agent’s trajectory can be divided into discrete behavioral states with corresponding unique reward function. Nevertheless all these algorithms assume linear reward functions and the Bayesian inference problem is intractable even for moderately sized finite-state IRL problems. Finally, the framework from Ashwood et al. (2022a) parametrizes the animal’s reward function as a time-varying linear combination of a small number of spatial reward maps with Gaussian random walk prior over weights, capturing continuous time-varying reward functions. Their approach pursue a related aim to ours, yet is limited to capturing continuous intra-episode variation of reward functions (as illustrated in Section 5.1), and difficult to adopt to other environments. Last but not least, most of the aforementioned algorithms are model-based, relying on a known transition dynamics of the environment, whereas in many scenarios, the environment model is unknown.
3 Background
3.1 Inverse Reinforcement Learning
Consider a Markov Decision Process (MDP): $\{{\mathcal{S}},{\mathcal{A}},T,{\mathcal{R}},\gamma\}$, where ${\mathcal{S}}$ and ${\mathcal{A}}$ denotes the state- and action-space; $T\colon{\mathcal{S}}\times{\mathcal{A}}\to\Delta({\mathcal{S}})$ is the state transition function ($\Delta$ denotes the probability simplex) with $T(s,a,s^{\prime})\coloneqq\Pr(s^{\prime}\mid s,a)$; ${\mathcal{R}}\colon{\mathcal{S}}\times{\mathcal{A}}\to{\mathbb{R}}$ defines the reward function, and $\gamma\in[0,1)$ denotes the discount factor. The problem of inverse reinforcement learning is formally defined as:
Problem 1 (IRL problem).
Given the demonstration space ${\mathcal{D}}\coloneqq{\{\xi_{i}\}}_{i=1}^{N}$ with $N$ trajectories, where each trajectory is a sequence of state-action pairs, $\xi_{i}\coloneqq\{(s_{1},a_{1}),(s_{2},a_{2}),\ldots\}$, the IRL problem consists of finding a reward function ${\mathcal{R}}$ that maximizes the LL between agent demonstrations and the (soft) optimal policy $\pi_{{\mathcal{R}}}$ under ${\mathcal{R}}$:
$\operatorname*{maximize}_{{\mathcal{R}}}\quad\sum_{i=1}^{N}\log(\Pr(\xi_{i}\mid\pi_{{\mathcal{R}}}))\text{\,.}$ (1)
3.2 Inverse Q-learning
The class of Inverse Q-learning algorithms (Kalweit et al., 2020) provides a precise yet notably time-efficient solution to Problem 1, compared to the popular Maximum Entropy IRL algorithm from Ziebart et al. (2008) and some of its variants. It assumes that the demonstrations are collected from an agent following a Boltzmann policy according to its unknown optimal value function $Q^{*}$:
$\pi^{E}(s,a)\coloneqq\frac{\exp(Q^{*}(s,a))}{\sum_{A\in{\mathcal{A}}}\exp(Q^{*}(s,A))}\Rightarrow Q^{*}(s,a)=Q^{*}(s,b)+\log(\pi^{E}(s,a))-\log(\pi^{E}(s,b))\text{\,,}$ (2)
for all actions $a\in{\mathcal{A}}$ and $b\in{\mathcal{A}}_{\bar{a}}$ where ${\mathcal{A}}_{\bar{a}}\coloneqq{\mathcal{A}}\backslash\{a\}$. Using the Bellman optimality equation in Equation 2, the immediate reward of action $a$ in state $s$ can be expressed by the immediate reward of some other action $b\in{\mathcal{A}}_{\bar{a}}$, the respective log-probabilities and future action-values:
$r(s,a)=\eta_{s}^{a}+\frac{1}{d_{{\mathcal{A}}}-1}\sum_{b\in{\mathcal{A}}_{\bar{a}}}\left[r(s,b)-\eta_{s}^{b}\right]\text{\,,}$ (3)
where $r(s,a)\in{\mathcal{R}}$ is the unknown reward function, $d_{{\mathcal{A}}}$ denotes the dimension of ${\mathcal{A}}$, and $\eta_{s}^{a}\coloneqq\log(\pi^{E}(s,a))-\gamma\sum_{s^{\prime}\in{\mathcal{S}}}T(s,a,s^{\prime})\max_{a^{\prime}\in{\mathcal{A}}}Q^{*}(s^{\prime},a^{\prime})$. The resulting system of linear equations can be solved with least squares, leading to the model-based Inverse Action-value Iteration (IAVI) algorithm, which solves the IRL problem analytically in closed-form. To relax the assumption of an existing transition model and action probabilities, IAVI was further extended to the sampling-based model-free Inverse Q-learning (IQL) algorithm (Kalweit et al., 2020). They showed that the Boltzmann distribution induced by the optimal action-value function on the learned reward from IAVI and IQL is equivalent to the arbitrary demonstrated behavior distribution.
4 Inverse Q-learning about Multiple Intentions
We first define the Multiple Intention Inverse Reinforcement Learning (MI-IRL) problem accordingly:
Problem 2 (MI-IRL problem).
Let ${\mathcal{Z}}\coloneqq{\{z_{k}\}}_{k=1}^{K}$ be a $K$-dimensional latent state space with each $z_{k}\in{\mathcal{Z}}$ corresponding to one intention, and let ${\mathcal{D}}\coloneqq{\{\xi_{i}\}}_{i=1}^{N}$ be $N$ trajectories demonstrated by an agent each under one of the latent states without labels. The MI-IRL problem consists of inferring the latent state labels and the corresponding reward functions ${\{{\mathcal{R}}_{k}\}}_{k=1}^{K}$ in ${\mathcal{D}}$ such that under the $k$-th latent state the agent (softly) optimizes ${\mathcal{R}}_{k}$.
We adopt the EM (Dempster et al., 1977) as a straightforward approach to attack Problem 2. Let $\Theta$ be the set of parameters to be inferred, and let ${\mathcal{Y}}\coloneqq{\{y_{i}\}}_{i=1}^{N}$ be the set of latent state labels for each trajectory $\xi\in{\mathcal{D}}$, where $y_{i}=k$ if trajectory $i$ came from under latent state $z_{k}$. At iteration $\tau$ of the EM process before $\Theta$ converges, the expected value of the likelihood function ${\mathcal{L}}$ of $\Theta$ will be maximized as in the following update equation:
$\Theta^{\tau+1}\coloneqq\operatorname*{arg\,max}_{\Theta}\sum_{{\mathcal{Y}}}{\mathcal{L}}(\Theta\mid{\mathcal{D}},{\mathcal{Y}})\Pr({\mathcal{Y}}\mid{\mathcal{D}},\Theta^{\tau})$ (4)
Noting that different latent state transition dynamics lead to respective parameter space $\Theta$ and specific implementations of Equation 4. In the following, we consider the latent state transition dynamics described with a generalized Bernoulli process (independent latent states) and a Markov process (Markovian interdependent latent states).
4.1 Clustering of Independent Latent States
We start from the simpler case where the occurrence of different intentions satisfies a generalized Bernoulli process. Let $\{\nu_{1},\ldots,\nu_{K}\mid\nu_{1}+\cdots+\nu_{K}=1\}$ be the set of prior probability corresponding to the occurrence of each latent state $z_{k}$, the set of parameters $\Theta$ to be inferred is then $\{\nu_{1},\ldots,\nu_{K};{\mathcal{R}}_{1},\ldots,{\mathcal{R}}_{K}\mid\nu_{1}+\cdots+\nu_{K}=1\}$. The optimal value for respective parameters in this parameter set at each EM iteration is provided by Theorem 1:
Theorem 1.
Given that the intention transition dynamics satisfies a generalized Bernoulli process, at iteration $\tau$, the EM update equation (Equation 4) for each parameter in the corresponding parameter set $\Theta=\{\nu_{1},\ldots,\nu_{K};{\mathcal{R}}_{1},\ldots,{\mathcal{R}}_{K}\mid\nu_{1}+\cdots+\nu_{K}=1\}$ is given by
$&\nu_{k}^{\tau+1}\coloneqq\frac{1}{N}\sum_{i=1}^{N}\zeta_{ik}^{\tau}\\ &{\mathcal{R}}_{k}^{\tau+1}\coloneqq\operatorname*{arg\,max}_{{\mathcal{R}}_{k}}\sum_{i=1}^{N}\zeta_{ik}^{\tau}\log(\Pr(\xi_{i}\mid\pi_{{\mathcal{R}}_{k}}))\text{\,,}\end{aligned}\right.$ (5)
for all $\nu,{\mathcal{R}}\in\Theta$, where $\zeta_{ik}\coloneqq\frac{1}{Z}\nu_{k}\prod_{(s,a)\in\xi_{i}}\pi_{{\mathcal{R}}_{k}}(s,a)$ is the probability that trajectory $i$ was demonstrated under latent state $z_{k}$ normalized by factor $Z$.
Proof.
See Appendix A.1. ∎
Noting that updating the reward function estimation ${\mathcal{R}}$ at each iteration according to Equation 5 is equivalent to solving Problem 1, except that each trajectory is weighted by a probability $\zeta$ during sampling. Thus combining the above EM approach for trajectory clustering with IAVI or IQL algorithms leads to the class of Latent Variable Inverse Q-learning (LV-IQL) algorithms (cf. Algorithm 1), solving the MI-IQL problem (Problem 2) when the occurrence of different latent states is independent.
4.2 Clustering of Markovian Interdependent Latent States
In addition to the generalized Bernoulli process, the Markov process is also considered an alternative for describing intention transition dynamics (Ashwood et al., 2022b; Le et al., 2023), where the occurrence of the next latent state corresponding to a specific intention is dependent and only dependent on the previous latent state. Under this assumption, given the agent demonstrations ${\mathcal{D}}$ consisting of a sequence of trajectories, the set of parameters to be inferred is then $\{\Pi,\Lambda;{\mathcal{R}}_{1},\ldots,{\mathcal{R}}_{K}\}$, where $\Pi\colon\Delta({\mathcal{Z}})$ and $\Lambda\colon{\mathcal{Z}}\to\Delta({\mathcal{Z}})$ ($\Delta$ is the probability simplex) denoting the latent state initial distribution probability and latent state transition matrix, respectively. The optimal value for respective parameters in the parameter set $\Theta$ at each EM iteration is provided by Theorem 2:
Theorem 2.
Given that the intention transition dynamics satisfies a Markov process, at iteration $\tau$, the EM update equation (Equation 4) for each parameter in the corresponding parameter set $\Theta=\{\Pi,\Lambda;{\mathcal{R}}_{1},\ldots,{\mathcal{R}}_{K}\}$ is given by [^1] Here we assume the index of trajectories in ${\mathcal{D}}$ starts from 0 instead of 1 for convenience but without losing generality.
$&\pi_{k}^{\tau+1}\coloneqq\Pr(y_{0}=k\mid{\mathcal{D}},\Theta^{\tau})\\ &\Lambda_{kl}^{\tau+1}\coloneqq\frac{\sum_{i=1}^{N}\Pr(y_{i-1}=k,y_{i}=l\mid{\mathcal{D}},\Theta^{\tau})}{\sum_{i=1}^{N}\Pr(y_{i-1}=k\mid{\mathcal{D}},\Theta^{\tau})}\\ &{\mathcal{R}}_{k}^{\tau+1}\coloneqq\operatorname*{arg\,max}_{{\mathcal{R}}_{k}}\sum_{i=0}^{N}\Pr(y_{i}=k\mid{\mathcal{D}},\Theta^{\tau})\log(\Pr(\xi_{i}\mid\pi_{{\mathcal{R}}_{k}}))\text{\,,}\end{aligned}\right.$ (6)
for all $\pi,\Lambda,{\mathcal{R}}\in\Theta$.
Proof.
See Appendix A.2. ∎
In practice, the Forward-Backward algorithm (Baum et al., 1970) can be used to address the probabilities in Equation 6, and IAVI or IQL then estimates the corresponding reward function independently for each latent state. This leads to the class of Latent Markov Variable Inverse Q-learning (LMV-IQL) algorithms (cf. Algorithm 2, details see also Appendix B.2), which solves Problem 2 when the latent state transition satisfies the Markov property.
5 Experiments
5.1 Application of LV-IQL to Simulated Behavior
Figure 1: Gridworld environment architecture with agent trajectories.
[…]
We compared between the performance of LV-IAVI, LV-IQL, IAVI, and IQL trained on the whole demonstration space. Two latent states were considered for both LV-IAVI and LV-IQL. As a measure of performance, we used the Expected Value Difference (EVD) metric (Levine et al., 2011). EVD is defined as the mean square error between the state-value under the true reward function for the expert policy and the state-value under the true reward for the optimal Boltzmann policy w.r.t. the learnt reward. It provides an estimation of the sub-optimality of the learnt policy under the true reward function. For LV-IAVI and LV-IQL, the inferred latent states with respective trajectory clusters were assigned to the best-fit ground truth intentions. Since IAVI and IQL assumed all trajectories were demonstrated under one intention, the EVD was analyzed twice on the ground truth reward for different intentions with the same learnt Boltzmann policy (Figure 2). The trajectory clusters learnt with LV-IAVI and LV-IQL are highly overlapped with the ground truth trajectory distribution. As a result, the learnt reward functions via LV-IAVI and LV-IQL match the respective ground truth reward functions exactly, while the single intention IAVI and IQL only resulted in a large EVD of $\sim 21$ for the ‘hungry’ intention and $\sim 38$ for the ‘thirsty’ intention, representing a mixed reward function for the two intentions. Similar results were found when we removed the punishment on intention irrelevant rewards (Appendix C.2). Noting that the DIRL algorithm (Ashwood et al., 2022a) is infeasible here as it assumes continuously time-varying rewards, which only addresses cases where the intention transition occurs after each action execution. However, in the above Gridworld experiment, each episode was conducted under one of two intentions, where the intention remains fixed within the episode, and the transition between intentions only occurs between episodes.
LV-IAVI LV-IQL IAVI IQL
EVD ‘Hungry’ $\bm{0.00\pm 0.00}$ $\bm{0.00\pm 0.00}$ $21.42\pm 0.00$ $21.38\pm 0.00$
‘Thirsty’ $\bm{0.00\pm 0.00}$ $\bm{0.00\pm 0.00}$ $38.00\pm 0.00$ $38.05\pm 0.00$
Figure 2: (Top) Visualization of the normalized ground truth and learnt state-value functions. Red lines indicate the ground truth trajectory distribution and the learnt trajectory clusters used to recover the reward function for respective intention. (All expert trajectories are shown for each figure of IAVI and IQL.) (Bottom) EVD for different approaches, Mean $\pm$ SE over 5 repeated runs.
5.2 Application of LMV-IAVI to Mice Navigating Trajectories
[…]
Figure 3: (A) The labyrinth environment. Blue line shows the optimal path from entrance to water port. Left, right, and reverse actions are represented with arrows while stay is denoted with cross. (B) Comparison of LMV-IAVI on test set trajectories to a random policy and DIRL, represented as LL. (C) BIC as a function of latent state numbers in LMV-IAVI. (D) Learnt policy (red arrows and crosses) in the environment and corresponding state occupancy (grey colormap) under different intentions. State occupancy was calculated by assigning each trajectory to the latent state with hightest posterior probability. Policies are shown only for states with non-zero occupancy. (E) Trajectories of latent state probabilities. Solid and shaded curves denote the Mean and SE. (F) Inferred latent state transition matrix from the best-fitting LMV-IAVI.
[…]
We demonstrate model comparison by first applying our method to trajectories from the water-restricted animals. The test set log-likelihood (LL) is similar for LMV-IAVI and DIRL under single intention ($K=1$). However, LMV-IAVI with $K>2$ substantially outperforms DIRL (Figure 3). Although the test LL continues to grow for larger $K$, the Bayesian information criterion (BIC) appears to increase (Figure 3). Thus LMV-IAVI with 2 latent states is considered for subsequent analysis. The learnt mice policy under latent state 1 (‘Tired’) displays a preference of moving out from the water port towards the maze entrance and stay, while the policy under latent state 2 (‘Thirsty’) guides the mice directly to the water port along the optimal track. Correspondingly, in the ‘Tired’ latent state, the highest state occupancy is noted at the entrance state, while under ‘Thirsty’, it is observed at the water port (Figure 3). To delve into the intention transition dynamics, we computed the posterior probability over mice’s latent state across all trajectories. The recovered average temporal latent state trajectories show a high probability of the ‘Thirsty’ latent state at the beginning but later on tailed off, as the ‘Tired’ latent state gradually became dominant (Figure 3). These findings demonstrate that LMV-IAVI not only excels the state-of-the-art in predicting mice labyrinth navigating behavior, but also provides distinct and interpretable reward functions. Similarly, our LMV-IAVI algorithm again outperforms DIRL when applied to the water-unrestricted animal trajectory dataset. Further details can be found at Appendix D.2.
5.3 Application of LMV-IQL to Mice Reversal-learning Behavior
[…]
Figure 4: (A) LL (Mean $\pm$ SE, 5-fold cross-validation) as a function of $\ell_{h}$ of single latent state LMV-IQL. (B) Change in test LL as a function of latent state numbers in LMV-IQL with $\ell_{h}=3$, relative to the fQ-learning model (labeled ‘F.’). Each trace represents a single mouse, averaged over cross-validation. Solid black indicates the mean across animals, and the dashed curve indicates the example mouse. (C) BIC as a function of latent state numbers in LMV-IQL with $\ell_{h}=3$. (D) Learnt mice policy represented with the probability of switch, win-stay, and lose-switch. Each grey curve denotes one mouse. (E) Average task performance and trajectories of latent state probabilities. Solid and shaded curves denote the Mean and SE. (F) Inferred latent state transition matrix from the best-fitting LMV-IQL for the example mouse. (G) Overall task performance (gray) and the performance under different latent states. (H) Relationship between the probability of the ‘Exploitation’ latent state, 5 trials before block switch and the mean probability of the ‘Exploration’ latent state, 5 trials after block switch, Mean $\pm$ SE.
We begin our application of LMV-IQL on the recorded mice behavior by selecting the hyper-parameter $\ell_{h}$. At this step, we only consider single latent state LMV-IQL (equivalent to IQL). We compared the LL on training and test sets of multiple IQL fitting with different $\ell_{h}$ (Figure 4). The LL on test sets shows a bell-shaped curve as $\ell_{h}$ increases, indicating an overfit on the training set when $\ell_{h}>3$. Noting that there is an abnormal drop on training set LL at large $\ell_{h}$s. This can be explained with the insufficient sampling given the fixed set of expert demonstrations, since the size of the state space ${\mathcal{S}}$ grows exponentially as the history length $\ell_{h}$ increases. The best test LL is achieved at $\ell_{h}=3$, which is selected for subsequent steps. Next, to determine the number of intentions $K$ under which mice demonstrated the trajectories, we fit multiple LMV-IQL with varying numbers of latent states. In this step, we additionally applied a forgetting Q-learning (fQ-learning) model (Beron et al., 2022), which has been widely recognized as a prominent forward behavioral model for the reversal-learning task. This was done using the same dataset, serving as a baseline for comparative analysis. We found that the multiple intention LMV-IQL fitting substantially outperformed the single intention models (Figure 4). Although the BIC w.r.t. different $K$ indicates that both $K=2$ and $K=3$ are reasonable values (Figure 4), we will focus subsequent analysis on the LMV-IQL with 3 latent states for biological interpretability. (More details about LMV-IQL fitting see also Appendix E.)
The inferred mice policies from LMV-IQL define how the subjects make decisions under three intentions (Figure 4). One of these policies, operating within latent state 1, displays a strong inclination toward adopting a ‘win-stay’ and ‘lose-switch’ strategy, which is the optimal policy in this deterministic reward bandit task. On the other hand, within latent state 2, the policy, referred to as the ‘Win-stay’ policy, exhibits a preference for exploitation when the previous trial was successful. However, following error trials, it employs a random action selection strategy, indicated by a $\sim 0.5$ probability of executing a ‘lose-switch’. Lastly, in latent state 3, a characteristic ‘Exploration’ policy emerges, where the subject consistently favors selecting the option opposite to the one chosen in the preceding trial, irrespective of whether they had won or lost in that particular instance. The recovered latent state trajectories in the example session reveals that the most probable latent state often exhibits a probability close to 1, indicating a high degree of confidence in discerning the subject’s intent based on the observed data (Figure 5). The ‘Exploration’ intention predominantly manifested at the onset of a block and endured for a relatively brief duration, in alignment with the learned latent state transition matrix (Figure 4).
Figure 5: Posterior latent state probabilities for an example session. Dots and triangles indicate mice behavior.
The significant values along the diagonal of the transition matrix within latent states 1 and 2, corresponding to ‘Exploitation’ and ‘Win-stay’, signify a heightened preference for persisting in the same latent state over multiple consecutive trials. Additionally, it becomes evident that error trials tend to coincide with the trials where the posterior probability of ‘Win-stay’ and ‘Exploration’ latent states reaches its zenith, corroborating the presence of suboptimal exploratory behavior associated with these two intentions. The average latent state transition trajectories across all blocks closely resembles those observed in the example session (Figure 4). As each block begin with the animal’s performance at a relatively low level, there is a decline in the posterior probability associated with the ‘Exploitation’ latent state, accompanied by an increase in the probabilities of the other two latent states associated with suboptimal exploratory strategies. Nonetheless, as the subjects’ performance steadily improves, the ‘Exploitation’ latent state progressively reasserts its dominance. Finally, to quantify latent state occupancies across all sessions, we assigned each trial to its most probable state. In contrast to the cohort’s general correct rate of $0.74\pm 0.02$, mice performed significantly better within the ‘Exploitation’ latent state, achieving a correctness rate of $0.86\pm 0.01$. In comparison, they attained lower correctness rates of $0.64\pm 0.03$ and $0.56\pm 0.06$ in the two alternative latent states (Figure 4). Furthermore, it’s worth noting that the mean posterior probability of the ‘Exploration’ latent state at the beginning of a new block shows a positive correlation with the average probability of the ‘Exploitation’ state at the end of the preceding block (Figure 4), suggesting that the ‘Exploration’ latent state appears to involve a deliberate, exploration-oriented action selection when mice are highly engaged and possess a good understanding of the environment.
6 Conclusion
In this study, we introduce a novel class of Latent (Markov) Variable Inverse Q-learning (L(M)V-IQL) algorithms for characterizing animal behavior during complex decision-making tasks. We extend the class of IQL algorithms (Kalweit et al., 2020) to learn multiple discrete reward functions from demonstrations. Specifically, we address the two most prevalent types of intention transition dynamics: the generalized Bernoulli process and the Markov process, under both model-based and model-free contexts. To validate our framework and compare with the state-of-the-art, we conduct experiments on simulated and real animal behavior data. Our approaches demonstrate a substantial improvement in behavior prediction compared to DIRL (Ashwood et al., 2022a) on mice navigation trajectories (indicated by the LL on held-out trajectories), without losing interpretability of the learnt reward functions. Moreover, our method provides distinct and interpretable reward functions for the mice cohort engaged in the reversal-learning task, where the animals displayed a pattern of alternating between exploitation and exploration intentions, which could extend over several consecutive trials within a single session. The transitions between these intentions followed a typical block-correlated trajectory, wherein the mice were more likely to exhibit in exploratory behaviors at the start of a new block, particularly if they had been highly engaged in the task in the preceding block.
A compelling avenue for future research lies in extending our framework to involve function approximations, which would enable the learning of a low-dimensional embedding of each state in the environment via e.g. a deep neural network. Such extension would allow us to scale our approach to high-dimensional or continuous state spaces, while also enabling the generalization across states. Another promising direction would be to extend the fixed intention transition probabilities with e.g. a generalized linear model, to incorporate the identification of potential external factors that influence intention transition dynamics.
Acknowledgments
This work has been funded as part of BrainLinks-BrainTools, which is funded by the Federal Ministry of Economics, Science and Arts of Baden-Württemberg within the sustainability programme for projects of the Excellence Initiative II; as well as the Bernstein Award 2012, the Research Unit 5159 “Resolving Prefrontal Flexibility” (Grant DI 1908/11-1), and the Deutsche Forschungsgemeinschaft (Grants DI 1908/3-1, DI 1908/11-1, and DI 1908/6-1), all to I.D.
Appendices
Appendix A Proof of Theorems
A.1 Proof of Theorem 1
Proof.
Substitute the parameters $\Theta=\{\nu_{1},\ldots,\nu_{K};{\mathcal{R}}_{1},\ldots,{\mathcal{R}}_{K}\mid\nu_{1}+\cdots+\nu_{K}=1\}$ under independent latent state assumption into the EM update equation (Equation 4) and unroll:
$\Psi(\Theta,\Theta^{\tau})$ $\coloneqq\sum_{{\mathcal{Y}}}{\mathcal{L}}(\Theta\mid{\mathcal{D}},{\mathcal{Y}})\Pr({\mathcal{Y}}\mid{\mathcal{D}},\Theta^{\tau})$ (from Equation 4)
$=\sum_{{\mathcal{Y}}}\sum_{i=1}^{N}\log(\nu_{y_{i}}\Pr(\xi_{i}\mid\pi_{{\mathcal{R}}_{y_{i}}}))\prod_{i^{\prime}=1}^{N}\Pr(y_{i^{\prime}}\mid\xi_{i^{\prime}},\Theta^{\tau})$ (A.1)
$=\sum_{y_{1}}\cdots\sum_{y_{N}}\sum_{i=1}^{N}\sum_{k=1}^{K}\delta_{k=y_{i}}\log(\nu_{k}\Pr(\xi_{i}\mid\pi_{{\mathcal{R}}_{k}}))\prod_{i^{\prime}=1}^{N}\Pr(y_{i^{\prime}}\mid\xi_{i^{\prime}},\Theta^{\tau})$ (A.2)
$=\sum_{k=1}^{K}\sum_{i=1}^{N}\log(\nu_{k}\Pr(\xi_{i}\mid\pi_{{\mathcal{R}}_{k}}))\sum_{y_{1}}\cdots\sum_{y_{N}}\delta_{k=y_{i}}\prod_{i^{\prime}=1}^{N}\Pr(y_{i^{\prime}}\mid\xi_{i^{\prime}},\Theta^{\tau})$ (A.3)
$=\sum_{k=1}^{K}\sum_{i=1}^{N}\log(\nu_{k}\Pr(\xi_{i}\mid\pi_{{\mathcal{R}}_{k}}))\zeta_{ik}^{\tau}$ (by Equation 5)
$=\sum_{k=1}^{K}\sum_{i=1}^{N}\zeta_{ik}^{\tau}\log(\nu_{k})+\sum_{k=1}^{K}\sum_{i=1}^{N}\zeta_{ik}^{\tau}\log(\Pr(\xi_{i}\mid\pi_{{\mathcal{R}}_{k}}))\text{\,,}$ (A.4)
where $\delta$ denotes the Kronecker delta function. Equation A.4 indicates that $\nu_{k}$ and ${\mathcal{R}}_{k}$ are not interdependent, we can thus optimize them separately in the M-step of EM, leading to the second update equation in Equation 5 trivially. According to Gibbs’ inequality, the first term of Equation A.4 is maximized if and only if
$\nu_{k}^{\tau+1}\coloneqq\frac{1}{N}\sum_{i=1}^{N}\zeta_{ik}^{\tau}\text{\,,}$ (A.5)
for all $\nu\in\Theta$, proving the first update equation in Equation 5. ∎
A.2 Proof of Theorem 2
Proof.
Similar to the proof for Theorem 1, substitute the parameter set $\Theta=\{\Pi,\Lambda;{\mathcal{R}}_{1},\ldots,{\mathcal{R}}_{K}\}$ into the EM update equation (Equation 4) and unroll:
$\Psi(\Theta,\Theta^{\tau})$ $\coloneqq\sum_{{\mathcal{Y}}}{\mathcal{L}}(\Theta\mid{\mathcal{D}},{\mathcal{Y}})\Pr({\mathcal{Y}}\mid{\mathcal{D}},\Theta^{\tau})$ (from Equation 4)
$=\sum_{{\mathcal{Y}}}\log(\pi_{y_{0}}\Pr(\xi_{0}\mid\pi_{{\mathcal{R}}_{y_{0}}})\prod_{i=1}^{N}\Lambda_{y_{i-1}y_{i}}\Pr(\xi_{i}\mid\pi_{{\mathcal{R}}_{y_{i}}}))\Pr({\mathcal{Y}}\mid{\mathcal{D}},\Theta^{\tau})$ (A.6)
$=\sum_{{\mathcal{Y}}}\log(\pi_{y_{0}})\Pr({\mathcal{Y}}\mid{\mathcal{D}},\Theta^{\tau})$
$\quad+\sum_{{\mathcal{Y}}}\sum_{i=1}^{N}\log(\Lambda_{y_{i-1}y_{i}})\Pr({\mathcal{Y}}\mid{\mathcal{D}},\Theta^{\tau})$
$\quad+\sum_{{\mathcal{Y}}}\sum_{i=0}^{N}\log(\Pr(\xi_{i}\mid\pi_{{\mathcal{R}}_{y_{i}}}))\Pr({\mathcal{Y}}\mid{\mathcal{D}},\Theta^{\tau})$ (A.7)
$=\sum_{y_{1}}\cdots\sum_{y_{N}}\sum_{k=1}^{K}\delta_{k=y_{0}}\log(\pi_{k})\prod_{i^{\prime}=1}^{N}\Pr(y_{i^{\prime}}\mid\xi_{i^{\prime}},\Theta^{\tau})$
$\quad+\sum_{y_{1}}\cdots\sum_{y_{N}}\sum_{i=1}^{N}\sum_{k=1}^{K}\sum_{l=1}^{K}\delta_{k=y_{i-1},l=y_{i}}\log(\Lambda_{kl})\prod_{i^{\prime}=1}^{N}\Pr(y_{i^{\prime}}\mid\xi_{i^{\prime}},\Theta^{\tau})$
$\quad+\sum_{y_{1}}\cdots\sum_{y_{N}}\sum_{i=0}^{N}\sum_{k=1}^{K}\delta_{k=y_{i}}\log(\Pr(\xi_{i}\mid\pi_{{\mathcal{R}}_{k}}))\prod_{i^{\prime}=1}^{N}\Pr(y_{i^{\prime}}\mid\xi_{i^{\prime}},\Theta^{\tau})$ (A.8)
$=\sum_{k=1}^{K}\Pr(y_{0}=k\mid{\mathcal{D}},\Theta^{\tau})\log(\pi_{k})$
$\quad+\sum_{k=1}^{K}\sum_{l=1}^{K}\sum_{i=1}^{N}\Pr(y_{i-1}=k,y_{i}=l\mid{\mathcal{D}},\Theta^{\tau})\log(\Lambda_{kl})$
$\quad+\sum_{k=1}^{K}\sum_{i=0}^{N}\Pr(y_{i}=k\mid{\mathcal{D}},\Theta^{\tau})\log(\Pr(\xi_{i}\mid\pi_{{\mathcal{R}}_{k}}))\text{\,,}$ (A.9)
where $\delta$ denotes the Kronecker delta function. Since $\pi_{k}$, $\Lambda_{kl}$ and ${\mathcal{R}}_{k}$ are not interdependent, we can thus maximize the respective term separately, resulting in Equation 6. ∎
Remark 1.
In a more practical case where the agent demonstration space has multiple trajectory sequences, Theorem 2 can also be generalized and proved in the same manner.
Remark 2.
All $\xi\in{\mathcal{D}}$ above are assumed to be the trajectory for a whole episode. In some special cases where it is assumed that the latent state transition happens after each action execution, instead of per episode, Theorem 2 can also be applied by regarding each episode as a trajectory sequence with each trajectory consists of only one action execution.
Appendix B Algorithms
B.1 Latent Variable Inverse Q-learning
The pseudo code for LV-IQL can be found at Algorithm 1.
Input: agent demonstrations ${\mathcal{D}}$, latent space dimension $K$
1 initialize $\Theta\coloneqq\{\nu_{1},\ldots,\nu_{K};{\mathcal{R}}_{1},\ldots,{\mathcal{R}}_{K}\mid\nu_{1}+\cdots+\nu_{K}=1\}$ repeat
2       E-step
3             foreach $\xi_{i}\in{\mathcal{D}}$ do
4                   forall $k$ do
5                         $\zeta_{ik}\leftarrow\prod_{(s,a)\in\xi_{i}}\pi_{{\mathcal{R}}_{k}}(s,a)\nu_{k}/Z$
6
7
8
9      M-step
10             forall $k$ do
11                   $\nu_{k}\leftarrow\sum_{i=1}^{N}\zeta_{ik}/N$ compute ${\mathcal{R}}_{k}$ via IAVI or IQL on ${\mathcal{D}}$ with weight $\zeta_{ik}$ on trajectory $\xi_{i}$
12
13
14until convergence;
Output: $\Theta$
Algorithm 1 Latent Variable Inverse Q-learning (LV-IQL)
B.2 Latent Markov Variable Inverse Q-learning
To implement the LMV-IQL algorithm, let the forward probability $\mathfrak{a}_{ik}$ be the posterior probability of the observed agent demonstrations up until trajectory $i$ and the latent state under which the $i$-th trajectory was demonstrated is $z_{k}$:
$\mathfrak{a}_{ik}$ $\coloneqq\Pr({\mathcal{D}}_{0:i},z_{i}=k\mid\Theta)$ (B.1)
=$&\Pi_{k}\Pr(\xi_{0}\mid z_{0}=k,\Theta),&&i=0\\ &\sum_{j=1}^{K}\mathfrak{a}_{(i-1)j}\Lambda_{jk}\Pr(\xi_{i}\mid z_{i}=k,\Theta),&&i\neq 0\text{\,,}\end{aligned}\right.$
and the backward probability $\mathfrak{b}_{ik}$ be the posterior probability of the demonstrations after trajectory $i$:
$\mathfrak{b}_{ik}$ $\coloneqq\Pr({\mathcal{D}}_{i+1:N}\mid z_{i}=k,\Theta)$ (B.2)
=$&\sum_{j=1}^{K}\mathfrak{b}_{(i+1)j}\Lambda_{kj}\Pr(\xi_{i+1}\mid z_{i+1}=j,\Theta),&&i\neq N\\ &1,&&i=N\text{\,.}\end{aligned}\right.$
The posterior probability that trajectory $i$ was demonstrated under latent state $z_{k}$ is then denoted as:
$\mathfrak{g}_{ik}$ $\coloneqq\Pr(y_{i}=k\mid{\mathcal{D}},\Theta)$ (B.3)
$=\frac{\mathfrak{a}_{ik}\mathfrak{b}_{ik}}{\sum_{j=1}^{K}\mathfrak{a}_{ij}\mathfrak{b}_{ij}}\text{\,,}$
and the posterior probability that trajectory $i-1$ was demonstrated under latent state $z_{k}$ and concomitantly trajectory $i$ was demonstrated under latent state $z_{l}$ is:
$\mathfrak{x}_{ikl}$ $\coloneqq\Pr(y_{i-1}=k,y_{i}=l\mid{\mathcal{D}},\Theta)$ (B.4)
$=\frac{\mathfrak{a}_{(i-1)k}\Lambda_{kl}\Pr(\xi_{i}\mid z_{i}=l,\Theta)\mathfrak{b}_{il}}{\sum_{u=1}^{K}\sum_{v=1}^{K}\mathfrak{a}_{(i-1)u}\Lambda_{uv}\Pr(\xi_{i}\mid z_{i}=v,\Theta)\mathfrak{b}_{iv}}\text{\,.}$
Thus the update equation in Equation 6 is equivalent to
$&\Pi_{k}^{\tau+1}\coloneqq\mathfrak{g}_{0k}^{\tau}\\ &\Lambda_{kl}^{\tau+1}\coloneqq\frac{\sum_{i=1}^{N}\mathfrak{x}_{ikl}^{\tau}}{\sum_{i=0}^{N-1}\mathfrak{g}_{ik}^{\tau}}\\ &{\mathcal{R}}_{k}^{\tau+1}\coloneqq\operatorname*{arg\,max}_{{\mathcal{R}}_{k}}\sum_{i=0}^{N}\mathfrak{g}_{ik}^{\tau}\log(\Pr(\xi_{i}\mid\Pi_{{\mathcal{R}}_{k}}))\text{\,.}\end{aligned}\right.$ (B.5)
Combining Equation B.5 and the class of IQL algorithms leads to LMV-IQL (Algorithm 2).
Input: expert demonstrations ${\mathcal{D}}$, latent space dimension $K$
1 initialize $\Theta\coloneqq\{\Pi,\Lambda;{\mathcal{R}}_{1},\ldots,{\mathcal{R}}_{K}\}$ repeat
2       E-step
3             calculate $\mathfrak{g}$ and $\mathfrak{x}$ according to Equations B.1B.4
4
5      M-step
6             forall $k$ do
7                   $\Pi_{k}\leftarrow\mathfrak{g}_{0k}$ forall $l$ do
8                         $\Lambda_{kl}\leftarrow\sum_{i=1}^{N}\mathfrak{x}_{ikl}/\sum_{i=0}^{N-1}\mathfrak{g}_{ik}$
9                  compute ${\mathcal{R}}_{k}$ via IAVI or IQL on ${\mathcal{D}}$ with weight $\mathfrak{g}_{ik}$ on trajectory $\xi_{i}$
10
11
12until convergence;
Output: $\Theta$
Algorithm 2 Latent Markov Variable Inverse Q-learning (LMV-IQL)
Appendix C Further Details and Additional Results on the Simulated Gridworld Behavior Dataset
C.1 The Gridworld Dataset and Model Training
The simulated agent demonstration space from the Gridworld environment consisted of 512 trajectories with each having a length of 64 movements. The discount factor was set to be $\gamma=0.99$. All evaluated algorithms were trained for 5 repeated runs on the whole demonstration space until convergence (difference of learnt reward function and the posterior probability of intentions for each trajectory $<10^{-3}$ between iterations).
C.2 Additional Results on Gridworld
LV-IAVI LV-IQL IAVI IQL
EVD ‘Hungry’ $\bm{0.00\pm 0.00}$ $\bm{1.00\pm 0.90}$ $3.03\pm 0.00$ $3.02\pm 0.00$
‘Thirsty’ $\bm{0.00\pm 0.00}$ $\bm{2.00\pm 1.79}$ $7.98\pm 0.00$ $7.98\pm 0.00$
Figure 6: (Top) Visualization of the normalized ground truth and learnt state-value functions. Red lines indicate the ground truth trajectory distribution and the learnt trajectory clusters used to recover the reward function for respective intention. (All expert trajectories are shown for each figure of IAVI and IQL.) (Bottom) EVD for different approaches, Mean $\pm$ SE over 5 repeated runs.
We also performed analysis under the environment set up where the intention irrelevant punishments were removed, i.e. replacing the $-1$ reward on the type of reward irrelevant to the intentions with 00. In this environment, there is an increased overlapping between some of the demonstrated trajectories under different intentions (Figure 6). However, LV-IAVI and LV-IQL still outperform the single intention algorithm IAVI and IQL in trajectory clustering and recovering corresponding expert reward functions.
Appendix D Further Details and Additional Results on the Evaluation of Mice Navigation Trajectories
D.1 Labyrinth Navigation Task and Model Training
In the navigation task from Rosenberg et al. (2021), two cohorts of 10 mice moved freely in dark through the labyrinth over the course of 7 hours. For comparability with the result from Ashwood et al. (2022a), we obtained their pre-processed mouse trajectories for water-restricted and water-unrestricted animals from https://github.com/97aditi/dynamic_irl [^2] The original recorded animal trajectories from Rosenberg et al. (2021) are provided with MIT open source license at https://github.com/markusmeister/Rosenberg-2021-Repository.. For the pre-processing, Ashwood et al. (2022a) used a clustering algorithm (based on DBSCAN (Ester et al., 1996)) for aligning trajectories across animals and bouts to reduce variability. After the pre-processing, they obtained 200 trajectories from the water-restricted animals and 207 trajectories from the water-unrestricted animals. 20% of trajectories from each cohort were held out as a test set.
To compare the performance of LMV-IAVI and DIRL in this environment, we used the source code provided by Ashwood et al. (2022a) to train DIRL on the animal trajectory dataset. All LMV-IAVI algorithms were trained for 10 repeated runs with different initializations, and the results from the initializations with hightest test set LL was selected for analysis. The initial latent state distribution $\Pi$ was initialized with a uniform distribution on the latent state space ${\mathcal{Z}}$ as $\Pi\coloneqq{\mathcal{U}}({\mathcal{Z}})$, and the latent state transition matrix $\Lambda$ was initialized as: $\Lambda\coloneqq 0.95\times I+{\mathcal{N}}(0,0.05\times I)$, where ${\mathcal{N}}$ denotes the normal distribution and $I\colon{\mathcal{Z}}\times{\mathcal{Z}}\to{\mathbb{R}}$ is the identity matrix. This initial $\Lambda$ was then normalized so that each row added up to 1. The discount factor was set to be $\gamma=0.99$.
D.2 Additional Results for Water-unrestricted Mice
[…]
Figure 7: (A) Comparison of LMV-IAVI on test set trajectories to a random policy and DIRL, represented as LL. (B) Learnt policy (red arrows) under different intentions and corresponding state occupancy (grey colormap) in the environment. State occupancy was calculated by assigning each trajectory to the latent state with hightest posterior probability. Policies are shown only for some states. (C) Trajectories of latent state probabilities. olid and shaded curves denote the Mean and SE. (D) Inferred latent state transition matrix from the best-fitting LMV-IAVI.
In contrast to the outcomes from the water-restricted animal dataset (Figure 3), LMV-IAVI demonstrates a higher test LL, even when considering a single intention. As the number of latent states (associated with DIRL’s goal maps) grows, the test LL of LMV-IAVI increases, while the test LL of DIRL remains constant (Figure 7). Focus on the two latent states LMV-IAVI, the inferred policy under two intentions exhibits ‘Exploring’ and ‘Tired’ behavior. The policy under ‘Exploring’ tends to encourage the animal lingering in the labyrinth, whereas the policy under ‘Tired’ latent state steers the animal back to the maze entrance. Correspondingly, the posterior probability of ‘Exploring’ initially dominates at the session’s beginning but is generally surpassed by the ‘Tired’ latent state over time.
Appendix E Further Details on the Evaluation of Mice Reversal-learning Behavior
The behavior data was collected from a cohort of mice consisted of 9 mice in total. Behavior recordings for each mice were repeated for at least 7 independent sessions with an average of $\sim 87$ trials per session.
We employed a multi-stage fitting procedure (Algorithm 3) to select hyper-parameters and to allow us to fit LMV-IQL individually to each animal. In the first stage, we concatenated the data from all animals in a single dataset together. We then performed multiple IQL (single latent state LMV-IQL) with different history truncation length $\ell_{h}\in\{1,\ldots,5\}$ on the concatenated data. Out of the 5 different values, We chose the $\ell_{h}$ that resulted in the best test set LL for subsequent stages. In the second stage, we run multiple LMV-IQL with different number of latent states $K\in\{2,\ldots,5\}$ again to the concatenated dataset to obtain a global fit. The initial latent state distribution $\Pi$ was initialized with a uniform distribution on the latent state space ${\mathcal{Z}}$ as $\Pi\coloneqq{\mathcal{U}}({\mathcal{Z}})$, and the latent state transition matrix $\Lambda$ was initialized as: $\Lambda\coloneqq 0.95\times I+{\mathcal{N}}(0,0.05\times I)$, where ${\mathcal{N}}$ denotes the normal distribution and $I\colon{\mathcal{Z}}\times{\mathcal{Z}}\to{\mathbb{R}}$ is the identity matrix. This initial $\Lambda$ was then normalized so that each row added up to 1. The reward and action-value function was initialized as $r(s,a)\coloneqq{\mathcal{N}}(0,0.2)$ and $Q(s,a)\coloneqq{\mathcal{N}}(0,5)$ for all $s\in{\mathcal{S}}$ and $a\in{\mathcal{A}}$. All discount factors were set to be $\gamma=0.99$. Since Algorithm 2 is not guaranteed to converge to the global optimum (Salakhutdinov et al., 2003), we performed 10 different initializations for each value of $K$. Out of the 10 initializations, we chose the parameters that resulted in the best training set LL for subsequent stages. In the last stage of the fitting procedure, we wanted to obtain a respective but aligned LMV-IQL fit for each animal, so we initialized the parameters for each animal with the best global fit parameters from all animals together, omitting the necessity to permute the retrieved latent states from each animal so as to map semantically similar intentions to one another. Algorithm 3 shows the pseudo-code for the whole procedure. A 5-fold cross-validation was used to split the training and test dataset [^3] Here we considered to hold out entire sessions of behavior for assessing test set performance. That is, the training and test set consisted of 80% and 20% of recorded sessions of each mouse, respectively., and Algorithm 3 was fit on each cross-validation fold independently.
Fit IQL globally:
1 foreach $\ell_{h}\in\{1,\ldots,5\}$ do
2       run IQL on the concatenated data from all animals until convergence
select best $\ell_{h}$ with largest test set LL Fit LMV-IQL globally:
3 foreach $K\in\{2,\ldots,5\}$ do
4       foreach $i\in\{1,\ldots,10\}$ do
5             initialize LMV-IQL with $K$ latent states and random parameters run Algorithm 2 on the concatenated data from all animals until convergence
6
Fit separate LMV-IQL to each animal:
7 forall animals do
8       foreach $K\in\{1,\ldots,5\}$ do
9             initialize LMV-IQL with $K$ latent states using the best global fit parameters for this $K$ run Algorithm 2 until convergence
10
Algorithm 3 Fitting LMV-IQL on real mice behavior
fBwB4ngcAcByHJpPJl+fn54mD3Gg0NrquXxeLRQAAwzAYj8cwTZPwPH9/sVg8PXweDAauqqr2cDjEer1GJBLBZDJBs9mE4zjwfZ85lAGg2+06hmGgXq+j3+/DsixYlgVN03a9Xu8jgCNCyIegIAgx13Vfd7vdu+FweG8YRkjXdWy329+dTgeSJD3ieZ7RNO0VAXAPwDEAO5VKndi2fWrb9jWl9Esul6PZbDY9Go1OZ7PZ9z/lyuD3OozU2wAAAABJRU5ErkJggg==" alt="[LOGO]">

### v2

@@
-L(M)V-IQL: Multiple Intention Inverse Reinforcement Learning for Animal Behavior Characterization
+Multi-intention Inverse Q-learning for Interpretable Behavior Representation
 Abstract
-In advancing the understanding of decision-making processes, mathematical models, particularly Inverse Reinforcement Learning (IRL), have proven instrumental in reconstructing animal’s multiple intentions amidst complex behaviors.
-Given the recent development of a continuous-time multi-intention IRL framework, there has been persistent inquiry into inferring discrete time-varying reward functions with multiple intention IRL approaches.
-To tackle the challenge, we introduce the Latent (Markov) Variable Inverse Q-learning (L(M)V-IQL) algorithms, a novel IRL framework tailored for accommodating discrete intrinsic rewards.
-Leveraging an Expectation-Maximization approach, we cluster observed trajectories into distinct intentions and independently solve the IRL problem for each.
+In advancing the understanding of decision-making processes, Inverse Reinforcement Learning (IRL) have proven instrumental in reconstructing animal’s multiple intentions amidst complex behaviors.
+Given the recent development of a continuous-time multi-intention IRL framework, there has been persistent inquiry into inferring discrete time-varying rewards with IRL.
+To tackle the challenge, we introduce Latent (Markov) Variable Inverse Q-learning (L(M)V-IQL), a novel class of IRL algorthms tailored for accommodating discrete intrinsic reward functions.
+Leveraging an Expectation-Maximization approach, we cluster observed expert trajectories into distinct intentions and independently solve the IRL problem for each.
 Demonstrating the efficacy of L(M)V-IQL through simulated experiments and its application to different real mouse behavior datasets, our approach surpasses current benchmarks in animal behavior prediction, producing interpretable reward functions.
-This advancement holds promise for neuroscience and psychology, contributing to a deeper understanding of animal decision-making and uncovering underlying brain mechanisms.
+This advancement holds promise for neuroscience and cognitive science, contributing to a deeper understanding of decision-making and uncovering underlying brain mechanisms.
+Inverse Reinforcement Learning, Neuroscience, Decision-making
 1 Introduction
-The characterization of animal’s decision-making behavior stands as a fundamental objective within the field of behavioral neuroscience (Niv, 2009; Wilson & Collins, 2019).
-Prior research has formulated a variety of mathematical behavioral models across diverse tasks, encompassing generalized linear models and models based on reinforcement learning (Ashwood et al., 2022b; Beron et al., 2022).
-Such forward models facilitate the comprehension and comparison of decision-making strategies employed by both human and animal subjects.
-Additionally, they offer a low-dimensional behavioral representation suitable for regression analysis with neural activities (Hattori et al., 2019; Hamaguchi et al., 2022).
-Forward models require a empirically defined reward function that guides subjects optimizing their behavior during decision-making.
-However, defining a comprehensive and suitable reward function can pose challenges in complex behavioral tasks.
-Alyahyay et al.
-(2023) introduced a response-preparation task where subjects ought to hold a lever until a cue indicating the release signal.
-In this task, subjects can receive a binary extrinsic reward from the environment, whereas the intrinsic reward driving behavior such as hunger, thirst, engagement, associated with each timestamp is, however, not obvious to the experimenter.
-As another example, within a 127-node-labyrinth with a water port at the terminal, Rosenberg et al.
-(2021) observed that the navigation behavior of water-restricted mice is influenced not solely by the extrinsic water reward but also by intrinsic motivators, including their curiosity to explore the environment.
-Inverse reinforcement learning (IRL) (Ng et al., 2000; Arora & Doshi, 2021) is a popular approach to recover a reward function that induces the observed behavior, assuming that the demonstrator was (softly) maximizing its long-term return.
-Along with the significant successes of IRL in autonomous driving (Kalweit et al., 2020; Nasernejad et al., 2023), robotics (Kumar et al., 2023; Chen et al., 2023), and healthcare domains (Coronato et al., 2020; Chan & van der Schaar, 2021), it appears to be emerging as an valuable tool for constructing mathematical behavior models in neuroscience research, as exemplified by Yamaguchi et al.
-(2018), Kwon et al.
-(2020), and Alyahyay et al.
-(2023).
-Classic IRL methods seek to identify a single, fixed reward function for a specific scenario.
-In contrast, Ashwood et al.
-(2022a) suggested that animal’s goals can evolve over time due to factors like fatigue, satiation, and curiosity.
-Under this assumption, they proposed the Dynamic Inverse Reinforcement Learning (DIRL) framework, which parametrizes the animal’s reward function as a time-varying linear combination of a small number of spatial reward maps, which are referred to as “goal maps”.
-By positing the existence of multiple goal maps with time-varying weights, DIRL allows the instantaneous reward function to vary continuously in time.
-This innovative framework achieved state-of-the-art performance in characterizing animal behavior.
-Nevertheless, persistent demands have emerged regarding an IRL framework incorporating discrete time-varying reward functions, particularly following the proposal by Ashwood et al.
-(2022b) that natural behaviors can be represented through a Markov chain characterized by alternating between discrete intentions.
-To address this requirement, we propose the novel class of Latent (Markov) Variable Inverse Q-learning (L(M)V-IQL) algorithms, which extend the single reward Inverse Q-learning (IQL) framework from Kalweit et al.
-(2020) to solve IRL problems accounting for multiple intentions.
-We formulate an Expectation-Maximization (EM) approach to first cluster animal trajectories into multiple intentions, and then solving the IRL problem independently for each intention, respectively.
-We theoretically demonstrate that L(M)V-IQL can cover the most common two types of intention transition dynamics: generalized Bernoulli process and Markov process.
-Finally, we present the application of our framework in 1) a simulated Gridworld environment, 2) real mice navigation trajectories with known environment model from Rosenberg et al.
-(2021), and 3) real mice decision-making data from a dynamic two-armed bandit task with unknown environment model, showing that our methods outperform state-of-the-art in behavior prediction, demonstrate exceptional aptitude in capturing the intentions of animals, and provide interpretable reward functions corresponding to each identified intention.
-2 Related Work
 Table 1: Overview of different multiple intention IRL algorithms.
@@
 BNP-IRL (Surana & Srivastava, 2014) $\times$ linear unknown discrete
+G-EM-MLIRL (Nguyen et al., 2015) $\times$ linear known discrete
 Meta-AIRL (Gleave & Habryka, 2018) $\checkmark$ non-linear $\times$ $\times$
@@
 L(M)V-IQL (Ours) $\checkmark$ non-linear known discrete
-Various approaches have been introduced to address multiple intention inverse reinforcement learning problems (Table 1).
-Notably, several frameworks based on parametric (Babes et al., 2011; Likmeta et al., 2021), or Bayesian non-parametric (Choi & Kim, 2012; Bighashdel et al., 2021) approaches share a similar objective to ours, allowing for multiple agents with distinct reward functions.
+Characterizing decision-making behavior stands as a fundamental objective within the field of behavioral neuroscience (Niv, 2009; Wilson & Collins, 2019).
+Prior research has formulated a variety of mathematical behavioral models across diverse tasks (Ashwood et al., 2022b; Beron et al., 2022), encompassing generalized linear models and models based on reinforcement learning.
+These forward models facilitate the comprehension and comparison of decision-making strategies employed by both human and animal subjects.
+Additionally, they offer a low-dimensional behavioral representation suitable for regression analysis with neural activities (Hattori et al., 2019; Hamaguchi et al., 2022).
+Forward models require an empirically defined reward function that guides subjects optimizing their behavior during decision-making.
+However, defining a comprehensive and suitable reward function can pose challenges in complex behavioral tasks.
+Alyahyay et al.
+(2023) introduced a response-preparation task where subjects ought to hold a lever until a cue indicating the release signal.
+In this task, subjects can receive a binary extrinsic reward from the environment, whereas the intrinsic reward driving behavior such as hunger, thirst, engagement, associated with each timestamp is, however, not obvious to the experimenter.
+As another example, within a 127-node-labyrinth with a water port at the terminal, Rosenberg et al.
+(2021) observed that the navigation behavior of water-restricted mice is influenced not solely by the extrinsic water reward but also by intrinsic motivators, including their curiosity to explore the environment.
+Inverse reinforcement learning (IRL) (Ng et al., 2000; Arora & Doshi, 2021) is a popular approach to recover a reward function that induces the observed behavior, assuming that the demonstrator was maximizing its long-term return.
+Along with the significant successes of IRL in autonomous driving (Kalweit et al., 2020; Nasernejad et al., 2023), robotics (Kumar et al., 2023; Chen et al., 2023), and healthcare domains (Coronato et al., 2020; Chan & van der Schaar, 2021), it appears to be emerging as an valuable tool for constructing mathematical behavior models in neuroscience research, as exemplified by Yamaguchi et al.
+(2018), Kwon et al.
+(2020), and Alyahyay et al.
+(2023).
+Classic IRL methods seek to identify a single, fixed reward function for a specific scenario.
+In contrast, Ashwood et al.
+(2022a) suggested that animal’s goals can evolve over time due to factors like fatigue, satiation, and curiosity.
+Under this assumption, they proposed the Dynamic Inverse Reinforcement Learning (DIRL) framework, which parametrizes the animal’s reward function as a smoothed time-varying linear combination of a small number of spatial reward maps, which are referred to as ‘goal maps’.
+By positing the existence of multiple goal maps with time-varying weights, DIRL allows the instantaneous reward function to vary continuously in time.
+This innovative framework achieved state-of-the-art performance in animal behavior prediction.
+Nevertheless, persistent demands have emerged regarding an IRL framework incorporating discrete time-varying reward functions, particularly following the proposal by Ashwood et al.
+(2022b) that natural behaviors can be represented through a Markov chain characterized by alternating between discrete intentions.
+To address this requirement, we propose the novel class of Latent (Markov) Variable Inverse Q-learning (L(M)V-IQL) algorithms, which extend the fixed-reward Inverse Q-learning (IQL) framework from Kalweit et al.
+(2020) to solve IRL problems accounting for multiple intentions.
+We formulate an Expectation-Maximization (EM) approach to first cluster animal trajectories into multiple intentions, and then solving the IRL problem independently for each intention.
+We theoretically demonstrate that L(M)V-IQL can cover the most common two types of intention transition dynamics: generalized Bernoulli process and Markov process.
+Finally, we present experiments on the application of our framework in 1) a simulated Gridworld environment; 2) real mice navigation trajectories with known environment model from the 127-node-labyrinth task (Rosenberg et al., 2021), serving as a benchmark for comparing with the state-of-the-art algorithm DIRL  (Ashwood et al., 2022a); and 3) real mice decision-making data with unknown environment model from a dynamic two-armed bandit task (De La Crompe et al., 2023).
+Demonstrating superior performance in behavior prediction, our methods showcase exceptional proficiency in capturing animals’ intentions through interpretable reward functions derived solely from their trajectories, surpassing the state-of-the-art approaches.
+2 Related Work
+Various approaches have been introduced to address Multiple Intention Inverse Reinforcement Learning (MI-IRL) problems (Table 1).
+Notably, several frameworks based on parametric (Babes et al., 2011; Likmeta et al., 2021), or non-parametric (Choi & Kim, 2012; Bighashdel et al., 2021) approaches allow for learning from multiple agents with distinct reward functions.
 However, these frameworks do not accommodate single agents with time-varying rewards.
 Gleave & Habryka (2018) developed a meta adversarial learning method for multi-task IRL problems.
-While their framework demonstrated high-level performance in real-world applications, it sacrifices theoretical interpretability and heavily relies on exploiting similarities between reward functions across tasks.
-Built on the Bayesian IRL approach, algorithms from Dimitrakakis & Rothkopf (2012), Michini & How (2012), and Surana & Srivastava (2014) posits that an agent’s trajectory can be divided into discrete behavioral states with corresponding unique reward function.
-Nevertheless all these algorithms assume linear reward functions and the Bayesian inference problem is intractable even for moderately sized finite-state IRL problems.
-Finally, the framework from Ashwood et al.
-(2022a) parametrizes the animal’s reward function as a time-varying linear combination of a small number of spatial reward maps with Gaussian random walk prior over weights, capturing continuous time-varying reward functions.
-Their approach pursue a related aim to ours, yet is limited to capturing continuous intra-episode variation of reward functions (as illustrated in Section 5.1), and difficult to adopt to other environments.
-Last but not least, most of the aforementioned algorithms are model-based, relying on a known transition dynamics of the environment, whereas in many scenarios, the environment model is unknown.
+While their framework demonstrated high-level performance in real-world applications, it sacrifices interpretability and heavily relies on exploiting similarities between reward functions across tasks.
+In contrast, several approaches formulate MI-IRL problem as finding the maximum-likelihood partition of each trajectory, where each segment was generated from different locally consistent reward functions.
+To solve the problem, Dimitrakakis & Rothkopf (2012), Michini & How (2012), and Surana & Srivastava (2014) established Bayesian IRL frameworks, while Nguyen et al.
+(2015) proposed a probabilistic graphical model, generalizing the algorithm from Babes et al.
+(2011).
+The latter approach avoids the computationally intensive Bayesian inference problem, which is intractable even for moderately sized finite-state IRL problems.
+Nevertheless all these algorithms are confined to linearly parameterized reward functions, whereas our L(M)V-IQL can effectively learn non-linear rewards.
+The state-of-the-art framework, DIRL (Ashwood et al., 2022a), parametrizes the expert’s reward function as a time-varying linear combination of a small number of non-linear spatial reward maps with Gaussian random walk prior over weights, capturing continuous time-varying rewards.
+Their approach pursues a related aim to ours, yet has the following limitations: 1) DIRL can only capture intra-episode variation of reward functions, while L(M)V-IQL can learn both intra- and inter-episode varying rewards, as discussed in Section 5.1.
+2) According to Ashwood et al.
+(2022b), humans and animals may switch between multiple discrete strategies during perceptual decision-making.
+Such behavioral characteristics would be challenging for DIRL to capture, as it assumes a smooth Gaussian prior on the time-varying rewards.
+Since the objective function and problem solver of DIRL heavily rely on this prior (Ashwood et al., 2022a), switching to other less smoothed prior would not be trivial.
+This limitation strongly affects DIRL’s performance, as demonstrated in Section 5.2, where we compare L(M)V-IQL and DIRL on the same benchmark.
+3) While DIRL excels at predicting mice navigation behavior in a labyrinth, it faces challenges when adapting to diverse behavioral tasks.
+As with other parametric frameworks, L(M)V-IQL adopts a design choice that requires a prior input specifying the number of intentions.
+This deliberate decision enables precise customization and enhances the algorithm’s efficiency in scenarios where intention specification is clear and well-defined.
+Last but not the least, most of the aforementioned algorithms are model-based, relying on a known transition dynamics of the environment, whereas in many scenarios, the environment model is unknown.
+As an improvement, our algorithms can perform model-free learning, enabling their application in a wider range of environments.
 3 Background
 3.1 Inverse Reinforcement Learning
-Consider a Markov Decision Process (MDP): $\{{\mathcal{S}},{\mathcal{A}},T,{\mathcal{R}},\gamma\}$, where ${\mathcal{S}}$ and ${\mathcal{A}}$ denotes the state- and action-space; $T\colon{\mathcal{S}}\times{\mathcal{A}}\to\Delta({\mathcal{S}})$ is the state transition function ($\Delta$ denotes the probability simplex) with $T(s,a,s^{\prime})\coloneqq\Pr(s^{\prime}\mid s,a)$; ${\mathcal{R}}\colon{\mathcal{S}}\times{\mathcal{A}}\to{\mathbb{R}}$ defines the reward function, and $\gamma\in[0,1)$ denotes the discount factor.
-The problem of inverse reinforcement learning is formally defined as:
-Problem 1 (IRL problem).
-Given the demonstration space ${\mathcal{D}}\coloneqq{\{\xi_{i}\}}_{i=1}^{N}$ with $N$ trajectories, where each trajectory is a sequence of state-action pairs, $\xi_{i}\coloneqq\{(s_{1},a_{1}),(s_{2},a_{2}),\ldots\}$, the IRL problem consists of finding a reward function ${\mathcal{R}}$ that maximizes the LL between agent demonstrations and the (soft) optimal policy $\pi_{{\mathcal{R}}}$ under ${\mathcal{R}}$:
-$\operatorname*{maximize}_{{\mathcal{R}}}\quad\sum_{i=1}^{N}\log(\Pr(\xi_{i}\mid\pi_{{\mathcal{R}}}))\text{\,.}$ (1)
+Consider the following Markov Decision Process (MDP): $\{\mathcal{S},\mathcal{A},T,r,\gamma\}$, where $\mathcal{S}$ and $\mathcal{A}$ denotes the state-space and action-space, respectively; $T\colon\mathcal{S}\times\mathcal{A}\to\mathcal{P}(\mathcal{S})$ is the state transition function ($\mathcal{P}$ is the probability simplex) with $T(s,a,s^{\prime})\coloneqq\Pr(s^{\prime}\mid s,a)$; $r\colon\mathcal{S}\times\mathcal{A}\to\mathbb{R}$ defines the reward function, and $\gamma\in[0,1)$ denotes the discount factor.
+Additionally, $\pi(s,a)\coloneqq\Pr(a\mid s)$ is used to represent the policy according to which actions are selected in the MDP.
+A formal definition of inverse reinforcement learning problems is then:
+Problem 3.1 (IRL problem).
+Given the demonstration space $\mathcal{D}\coloneqq{\{\xi_{i}\}}_{i=1}^{N}$ including $N$ trajectories provided by an expert in an MDP, where each $\xi_{i}\coloneqq\{(s_{1},a_{1}),(s_{2},a_{2}),\ldots\}$ is a sequence of state-action pairs, the IRL problem consists of finding a reward function $r$ that maximizes the log-likelihood between expert demonstrations and the optimal policy $\pi_{r}$ under $r$:
+maximize (over $r$) $\sum_{i=1}^{N}\log(\Pr(\xi_{i}\mid\pi_{r}))\text{\,.}$ (1)
 3.2 Inverse Q-learning
-The class of Inverse Q-learning algorithms (Kalweit et al., 2020) provides a precise yet notably time-efficient solution to Problem 1, compared to the popular Maximum Entropy IRL algorithm from Ziebart et al.
+The class of Inverse Q-learning algorithms (Kalweit et al., 2020) provides a precise yet notably time-efficient solution to 3.1, compared to the popular Maximum Entropy IRL algorithm from Ziebart et al.
 (2008) and some of its variants.
 It assumes that the demonstrations are collected from an agent following a Boltzmann policy according to its unknown optimal value function $Q^{*}$:
-$\pi^{E}(s,a)\coloneqq\frac{\exp(Q^{*}(s,a))}{\sum_{A\in{\mathcal{A}}}\exp(Q^{*}(s,A))}\Rightarrow Q^{*}(s,a)=Q^{*}(s,b)+\log(\pi^{E}(s,a))-\log(\pi^{E}(s,b))\text{\,,}$ (2)
-for all actions $a\in{\mathcal{A}}$ and $b\in{\mathcal{A}}_{\bar{a}}$ where ${\mathcal{A}}_{\bar{a}}\coloneqq{\mathcal{A}}\backslash\{a\}$.
-Using the Bellman optimality equation in Equation 2, the immediate reward of action $a$ in state $s$ can be expressed by the immediate reward of some other action $b\in{\mathcal{A}}_{\bar{a}}$, the respective log-probabilities and future action-values:
+$\pi^{E}(s,a)\coloneqq\frac{\exp(Q^{*}(s,a))}{\sum_{A\in\mathcal{A}}\exp(Q^{*}(s,A))}\text{\,.}$ (2)
+Rearranging Equation 2 leads to $Q^{*}(s,a)=Q^{*}(s,b)+\log(\pi^{E}(s,a))-\log(\pi^{E}(s,b))$, for all actions $a\in\mathcal{A}$ and $b\in\mathcal{A}_{\bar{a}}$ where $\mathcal{A}_{\bar{a}}\coloneqq\mathcal{A}\backslash\{a\}$.
+Substituting the Bellman optimality equation, i.e.
+$Q^{*}(s,a)=r(s,a)+\gamma\sum_{s^{\prime}\in\mathcal{S}}T(s,a,s^{\prime})\max_{a^{\prime}\in\mathcal{A}}Q^{*}(s^{\prime},a^{\prime})$ (Howard, 1960), the immediate reward of action $a$ at state $s$ can be expressed by the immediate reward of some other action $b\in\mathcal{A}_{\bar{a}}$, the respective log-probabilities and future action-values:
 $r(s,a)=\eta_{s}^{a}+\frac{1}{d_{\mathcal{A}}-1}\sum_{b\in\mathcal{A}_{\bar{a}}}\left[r(s,b)-\eta_{s}^{b}\right]\text{\,,}$ (3)
-where $r(s,a)\in{\mathcal{R}}$ is the unknown reward function, $d_{{\mathcal{A}}}$ denotes the dimension of ${\mathcal{A}}$, and $\eta_{s}^{a}\coloneqq\log(\pi^{E}(s,a))-\gamma\sum_{s^{\prime}\in{\mathcal{S}}}T(s,a,s^{\prime})\max_{a^{\prime}\in{\mathcal{A}}}Q^{*}(s^{\prime},a^{\prime})$.
+where $r(s,\cdot)$ is the unknown reward function at state $s\in\mathcal{S}$, $d_{\mathcal{A}}$ denotes the dimension of $\mathcal{A}$, and
+$\eta_{s}^{a}\coloneqq\log(\pi^{E}(s,a))-\gamma\sum_{s^{\prime}\in\mathcal{S}}T(s,a,s^{\prime})\max_{a^{\prime}\in\mathcal{A}}Q^{*}(s^{\prime},a^{\prime})\text{\,.}$ (4)
 The resulting system of linear equations can be solved with least squares, leading to the model-based Inverse Action-value Iteration (IAVI) algorithm, which solves the IRL problem analytically in closed-form.
@@
 4 Inverse Q-learning about Multiple Intentions
-We first define the Multiple Intention Inverse Reinforcement Learning (MI-IRL) problem accordingly:
-Problem 2 (MI-IRL problem).
+We first provide a formal definition of MI-IRL problems accordingly:
+Problem 4.1 (MI-IRL problem).
 Let $\mathcal{Z}\coloneqq{\{z_{k}\}}_{k=1}^{K}$ be a $K$-dimensional latent state space with each $z_{k}\in\mathcal{Z}$ corresponding to one intention, and let $\mathcal{D}\coloneqq{\{\xi_{i}\}}_{i=1}^{N}$ be $N$ trajectories demonstrated by an agent each under one of the latent states without labels.
-The MI-IRL problem consists of inferring the latent state labels and the corresponding reward functions ${\{{\mathcal{R}}_{k}\}}_{k=1}^{K}$ in ${\mathcal{D}}$ such that under the $k$-th latent state the agent (softly) optimizes ${\mathcal{R}}_{k}$.
-We adopt the EM (Dempster et al., 1977) as a straightforward approach to attack Problem 2.
+The MI-IRL problem consists of inferring the latent state labels and the corresponding reward functions ${\{r_{k}\}}_{k=1}^{K}$ in $\mathcal{D}$ such that under the $k$-th latent state the agent (softly) optimizes $r_{k}$.
+We adopt the EM (Dempster et al., 1977) as a straightforward approach to attack 4.1.
 Let $\Theta$ be the set of parameters to be inferred, and let $\mathcal{Y}\coloneqq{\{y_{i}\}}_{i=1}^{N}$ be the set of latent state labels for each trajectory $\xi\in\mathcal{D}$, where $y_{i}=k$ if trajectory $i$ came from under latent state $z_{k}$.
 At iteration $\tau$ of the EM process before $\Theta$ converges, the expected value of the likelihood function $\mathcal{L}$ of $\Theta$ will be maximized as in the following update equation:
-$\Theta^{\tau+1}\coloneqq\operatorname*{arg\,max}_{\Theta}\sum_{{\mathcal{Y}}}{\mathcal{L}}(\Theta\mid{\mathcal{D}},{\mathcal{Y}})\Pr({\mathcal{Y}}\mid{\mathcal{D}},\Theta^{\tau})$ (4)
+$\Theta^{\tau+1}\coloneqq\operatorname*{argmax}_{\Theta}\sum_{\mathcal{Y}}\mathcal{L}(\Theta\mid\mathcal{D},\mathcal{Y})\Pr(\mathcal{Y}\mid\mathcal{D},\Theta^{\tau})\text{\,.}$ (5)
 Noting that different latent state transition dynamics lead to respective parameter space $\Theta$ and specific implementations of Equation 5.
@@
 We start from the simpler case where the occurrence of different intentions satisfies a generalized Bernoulli process.
-Let $\{\nu_{1},\ldots,\nu_{K}\mid\nu_{1}+\cdots+\nu_{K}=1\}$ be the set of prior probability corresponding to the occurrence of each latent state $z_{k}$, the set of parameters $\Theta$ to be inferred is then $\{\nu_{1},\ldots,\nu_{K};{\mathcal{R}}_{1},\ldots,{\mathcal{R}}_{K}\mid\nu_{1}+\cdots+\nu_{K}=1\}$.
-The optimal value for respective parameters in this parameter set at each EM iteration is provided by Theorem 1:
-Theorem 1.
-Given that the intention transition dynamics satisfies a generalized Bernoulli process, at iteration $\tau$, the EM update equation (Equation 4) for each parameter in the corresponding parameter set $\Theta=\{\nu_{1},\ldots,\nu_{K};{\mathcal{R}}_{1},\ldots,{\mathcal{R}}_{K}\mid\nu_{1}+\cdots+\nu_{K}=1\}$ is given by
-$&\nu_{k}^{\tau+1}\coloneqq\frac{1}{N}\sum_{i=1}^{N}\zeta_{ik}^{\tau}\\ &{\mathcal{R}}_{k}^{\tau+1}\coloneqq\operatorname*{arg\,max}_{{\mathcal{R}}_{k}}\sum_{i=1}^{N}\zeta_{ik}^{\tau}\log(\Pr(\xi_{i}\mid\pi_{{\mathcal{R}}_{k}}))\text{\,,}\end{aligned}\right.$ (5)
-for all $\nu,{\mathcal{R}}\in\Theta$, where $\zeta_{ik}\coloneqq\frac{1}{Z}\nu_{k}\prod_{(s,a)\in\xi_{i}}\pi_{{\mathcal{R}}_{k}}(s,a)$ is the probability that trajectory $i$ was demonstrated under latent state $z_{k}$ normalized by factor $Z$.
+Let $\{\nu_{1},\ldots,\nu_{K}\mid\nu_{1}+\cdots+\nu_{K}=1\}$ be the set of prior probability corresponding to the occurrence of each latent state $z_{k}$, the set of parameters $\Theta$ to be inferred is then $\{\nu_{1},\ldots,\nu_{K};r_{1},\ldots,r_{K}\mid\nu_{1}+\cdots+\nu_{K}=1\}$.
+The optimal value for respective parameters in this parameter set at each EM iteration is provided by Theorem 4.2.
+Theorem 4.2.
+Given that the intention transition dynamics satisfies a generalized Bernoulli process, at iteration $\tau$, the EM update equation (Equation 5) for each parameter in the corresponding parameter set $\Theta=\{\nu_{1},\ldots,\nu_{K};r_{1},\ldots,r_{K}\mid\nu_{1}+\cdots+\nu_{K}=1\}$ is given by
+$&\nu_{k}^{\tau+1}\coloneqq\frac{1}{N}\sum_{i=1}^{N}\zeta_{ik}^{\tau}\\ &r_{k}^{\tau+1}\coloneqq\operatorname*{argmax}_{r_{k}}\sum_{i=1}^{N}\zeta_{ik}^{\tau}\log(\Pr(\xi_{i}\mid\pi_{r_{k}}))\text{\,,}\end{aligned}\right.$ (6)
+for all $\nu,r\in\Theta$, where
+$\zeta_{ik}\coloneqq\frac{1}{Z}\nu_{k}\prod_{(s,a)\in\xi_{i}}\pi_{r_{k}}(s,a)$ (7)
+is the probability that trajectory $i$ was demonstrated under latent state $z_{k}$ normalized by factor $Z$.
 Proof.
-See Appendix A.1. ∎
-Noting that updating the reward function estimation ${\mathcal{R}}$ at each iteration according to Equation 5 is equivalent to solving Problem 1, except that each trajectory is weighted by a probability $\zeta$ during sampling.
+See Section A.1. ∎
+Noting that updating the reward function estimation $r$ at each iteration according to Equation 6 is equivalent to solving 3.1, except that each trajectory is weighted by a probability $\zeta$ during sampling.
 Thus combining the above EM approach for trajectory clustering with IAVI or IQL algorithms leads to the class of Latent Variable Inverse Q-learning (LV-IQL) algorithms (cf.
-Algorithm 1), solving the MI-IQL problem (Problem 2) when the occurrence of different latent states is independent.
+Algorithm 1), solving the MI-IQL problem (4.1) when the occurrence of different latent states is independent.
 4.2 Clustering of Markovian Interdependent Latent States
-In addition to the generalized Bernoulli process, the Markov process is also considered an alternative for describing intention transition dynamics (Ashwood et al., 2022b; Le et al., 2023), where the occurrence of the next latent state corresponding to a specific intention is dependent and only dependent on the previous latent state.
-Under this assumption, given the agent demonstrations ${\mathcal{D}}$ consisting of a sequence of trajectories, the set of parameters to be inferred is then $\{\Pi,\Lambda;{\mathcal{R}}_{1},\ldots,{\mathcal{R}}_{K}\}$, where $\Pi\colon\Delta({\mathcal{Z}})$ and $\Lambda\colon{\mathcal{Z}}\to\Delta({\mathcal{Z}})$ ($\Delta$ is the probability simplex) denoting the latent state initial distribution probability and latent state transition matrix, respectively.
-The optimal value for respective parameters in the parameter set $\Theta$ at each EM iteration is provided by Theorem 2:
-Theorem 2.
-Given that the intention transition dynamics satisfies a Markov process, at iteration $\tau$, the EM update equation (Equation 4) for each parameter in the corresponding parameter set $\Theta=\{\Pi,\Lambda;{\mathcal{R}}_{1},\ldots,{\mathcal{R}}_{K}\}$ is given by [^1] Here we assume the index of trajectories in ${\mathcal{D}}$ starts from 0 instead of 1 for convenience but without losing generality.
-$&\pi_{k}^{\tau+1}\coloneqq\Pr(y_{0}=k\mid{\mathcal{D}},\Theta^{\tau})\\ &\Lambda_{kl}^{\tau+1}\coloneqq\frac{\sum_{i=1}^{N}\Pr(y_{i-1}=k,y_{i}=l\mid{\mathcal{D}},\Theta^{\tau})}{\sum_{i=1}^{N}\Pr(y_{i-1}=k\mid{\mathcal{D}},\Theta^{\tau})}\\ &{\mathcal{R}}_{k}^{\tau+1}\coloneqq\operatorname*{arg\,max}_{{\mathcal{R}}_{k}}\sum_{i=0}^{N}\Pr(y_{i}=k\mid{\mathcal{D}},\Theta^{\tau})\log(\Pr(\xi_{i}\mid\pi_{{\mathcal{R}}_{k}}))\text{\,,}\end{aligned}\right.$ (6)
-for all $\pi,\Lambda,{\mathcal{R}}\in\Theta$.
+In addition to the generalized Bernoulli process, the Markov process is also considered an alternative for describing intention transition dynamics (Ashwood et al., 2022b; Le et al., 2023).
+Under this assumption, given the agent demonstrations $\mathcal{D}$ consisting of a sequence of trajectories, the set of parameters to be inferred is then $\{\Pi,\Lambda;r_{1},\ldots,r_{K}\}$, where $\Pi\colon\mathcal{P}(\mathcal{Z})$ and $\Lambda\colon\mathcal{Z}\to\mathcal{P}(\mathcal{Z})$ ($\mathcal{P}$ is the probability simplex) denoting the latent state initial distribution probability and latent state transition matrix, respectively.
+The optimal value for respective parameters in the parameter set $\Theta$ at each EM iteration is provided by Theorem 4.3.
+Theorem 4.3.
+Given that the intention transition dynamics satisfies a Markov process, at iteration $\tau$, the EM update equation (Equation 5) for each parameter in the corresponding parameter set $\Theta=\{\Pi,\Lambda;r_{1},\ldots,r_{K}\}$ is given by
+$&\Pi_{k}^{\tau+1}\coloneqq\Pr(y_{0}=k\mid\mathcal{D},\Theta^{\tau})\\ &\Lambda_{kl}^{\tau+1}\coloneqq\frac{\sum_{i=1}^{N}\Pr(y_{i-1}=k,y_{i}=l\mid\mathcal{D},\Theta^{\tau})}{\sum_{i=1}^{N}\Pr(y_{i-1}=k\mid\mathcal{D},\Theta^{\tau})}\\ &r_{k}^{\tau+1}\coloneqq\operatorname*{argmax}_{r_{k}}\sum_{i=0}^{N}\Pr(y_{i}=k\mid\mathcal{D},\Theta^{\tau})\\ &\qquad\qquad\times\log(\Pr(\xi_{i}\mid\pi_{r_{k}}))\text{\,,}\end{aligned}\right.$ (8)
+for all $\Pi,\Lambda,r\in\Theta$ [^1] Here we assume the index of trajectories in $\mathcal{D}$ starts from 0 instead of 1 for convenience without losing generality..
 Proof.
-See Appendix A.2. ∎
+See Section A.2. ∎
 In practice, the Forward-Backward algorithm (Baum et al., 1970) can be used to address the probabilities in Equation 8, and IAVI or IQL then estimates the corresponding reward function independently for each latent state.
 This leads to the class of Latent Markov Variable Inverse Q-learning (LMV-IQL) algorithms (cf.
-Algorithm 2, details see also Appendix B.2), which solves Problem 2 when the latent state transition satisfies the Markov property.
+Algorithm 2, details see also Section B.2), which solves 4.1 when the latent state transition satisfies the Markov property.
 5 Experiments
 5.1 Application of LV-IQL to Simulated Behavior
-Figure 1: Gridworld environment architecture with agent trajectories.
+Figure 1: Gridworld environment with demonstrated trajectories.
 […]
@@
 As a result, the learnt reward functions via LV-IAVI and LV-IQL match the respective ground truth reward functions exactly, while the single intention IAVI and IQL only resulted in a large EVD of $\sim 21$ for the ‘hungry’ intention and $\sim 38$ for the ‘thirsty’ intention, representing a mixed reward function for the two intentions.
-Similar results were found when we removed the punishment on intention irrelevant rewards (Appendix C.2).
+Similar results were found when we removed the punishment on intention irrelevant rewards (Section C.2).
 Noting that the DIRL algorithm (Ashwood et al., 2022a) is infeasible here as it assumes continuously time-varying rewards, which only addresses cases where the intention transition occurs after each action execution.
@@
 Similarly, our LMV-IAVI algorithm again outperforms DIRL when applied to the water-unrestricted animal trajectory dataset.
-Further details can be found at Appendix D.2.
+Further details can be found at Section D.2.
 5.3 Application of LMV-IQL to Mice Reversal-learning Behavior
-[…]
 […]
@@
 (H) Relationship between the probability of the ‘Exploitation’ latent state, 5 trials before block switch and the mean probability of the ‘Exploration’ latent state, 5 trials after block switch, Mean $\pm$ SE.
+[…]
 We begin our application of LMV-IQL on the recorded mice behavior by selecting the hyper-parameter $\ell_{h}$.
@@
 On the other hand, within latent state 2, the policy, referred to as the ‘Win-stay’ policy, exhibits a preference for exploitation when the previous trial was successful.
+Figure 5: Posterior latent state probabilities for an example session. Dots and triangles indicate mice behavior.
 However, following error trials, it employs a random action selection strategy, indicated by a $\sim 0.5$ probability of executing a ‘lose-switch’.
@@
 The ‘Exploration’ intention predominantly manifested at the onset of a block and endured for a relatively brief duration, in alignment with the learned latent state transition matrix (Figure 4).
-Figure 5: Posterior latent state probabilities for an example session. Dots and triangles indicate mice behavior.
 The significant values along the diagonal of the transition matrix within latent states 1 and 2, corresponding to ‘Exploitation’ and ‘Win-stay’, signify a heightened preference for persisting in the same latent state over multiple consecutive trials.
@@
 Another promising direction would be to extend the fixed intention transition probabilities with e.g.
-a generalized linear model, to incorporate the identification of potential external factors that influence intention transition dynamics.
-Acknowledgments
+a generalized linear model (Nguyen et al., 2015), to incorporate the identification of potential external factors that influence intention transition dynamics.
+Finally, we also plan to extend our L(M)V-IQL algorithm to accommodate an unknown number of reward functions (Michini & How, 2012; Surana & Srivastava, 2014).
+Broader Impact
+Our work constitutes an advancement in IRL methods that can be used in behavioral neuroscience and cognitive science researches.
+Importantly, our work can be adapted to multiple behavioral experiment paradigms with minor modifications and enables the implementation by non-expert users.
+This is expected to boost productivity and tighten interdisciplinary research collaborations.
+On the other hand, we believe that our work also provides future directions in machine learning and robotics, such as the development of advanced reinforcement learning algorithms.
+By gaining insights into the intrinsic reward function driving animal behavior, we anticipate that our framework could serve as a valuable resource for formulating improved reward functions when training artificial agents or robots to perform challenging tasks.
+This can potentially accelerate existing trends for automation in industries.
+Acknowledgements
 This work has been funded as part of BrainLinks-BrainTools, which is funded by the Federal Ministry of Economics, Science and Arts of Baden-Württemberg within the sustainability programme for projects of the Excellence Initiative II; as well as the Bernstein Award 2012, the Research Unit 5159 “Resolving Prefrontal Flexibility” (Grant DI 1908/11-1), and the Deutsche Forschungsgemeinschaft (Grants DI 1908/3-1, DI 1908/11-1, and DI 1908/6-1), all to I.D.
@@
 Appendix A Proof of Theorems
-A.1 Proof of Theorem 1
+A.1 Proof of Theorem 4.2
 Proof.
-Substitute the parameters $\Theta=\{\nu_{1},\ldots,\nu_{K};{\mathcal{R}}_{1},\ldots,{\mathcal{R}}_{K}\mid\nu_{1}+\cdots+\nu_{K}=1\}$ under independent latent state assumption into the EM update equation (Equation 4) and unroll:
+Substitute the parameters $\Theta=\{\nu_{1},\ldots,\nu_{K};r_{1},\ldots,r_{K}\mid\nu_{1}+\cdots+\nu_{K}=1\}$ under independent latent state assumption into the EM update equation (Equation 5) and unroll:
 $\Psi(\Theta,\Theta^{\tau})$ $\coloneqq\sum_{\mathcal{Y}}\mathcal{L}(\Theta\mid\mathcal{D},\mathcal{Y})\Pr(\mathcal{Y}\mid\mathcal{D},\Theta^{\tau})$ (from Equation 5)
-$=\sum_{{\mathcal{Y}}}\sum_{i=1}^{N}\log(\nu_{y_{i}}\Pr(\xi_{i}\mid\pi_{{\mathcal{R}}_{y_{i}}}))\prod_{i^{\prime}=1}^{N}\Pr(y_{i^{\prime}}\mid\xi_{i^{\prime}},\Theta^{\tau})$ (A.1)
-$=\sum_{y_{1}}\cdots\sum_{y_{N}}\sum_{i=1}^{N}\sum_{k=1}^{K}\delta_{k=y_{i}}\log(\nu_{k}\Pr(\xi_{i}\mid\pi_{{\mathcal{R}}_{k}}))\prod_{i^{\prime}=1}^{N}\Pr(y_{i^{\prime}}\mid\xi_{i^{\prime}},\Theta^{\tau})$ (A.2)
-$=\sum_{k=1}^{K}\sum_{i=1}^{N}\log(\nu_{k}\Pr(\xi_{i}\mid\pi_{{\mathcal{R}}_{k}}))\sum_{y_{1}}\cdots\sum_{y_{N}}\delta_{k=y_{i}}\prod_{i^{\prime}=1}^{N}\Pr(y_{i^{\prime}}\mid\xi_{i^{\prime}},\Theta^{\tau})$ (A.3)
-$=\sum_{k=1}^{K}\sum_{i=1}^{N}\log(\nu_{k}\Pr(\xi_{i}\mid\pi_{{\mathcal{R}}_{k}}))\zeta_{ik}^{\tau}$ (by Equation 5)
-$=\sum_{k=1}^{K}\sum_{i=1}^{N}\zeta_{ik}^{\tau}\log(\nu_{k})+\sum_{k=1}^{K}\sum_{i=1}^{N}\zeta_{ik}^{\tau}\log(\Pr(\xi_{i}\mid\pi_{{\mathcal{R}}_{k}}))\text{\,,}$ (A.4)
+$=\sum_{\mathcal{Y}}\sum_{i=1}^{N}\log(\nu_{y_{i}}\Pr(\xi_{i}\mid\pi_{r_{y_{i}}}))\prod_{i^{\prime}=1}^{N}\Pr(y_{i^{\prime}}\mid\xi_{i^{\prime}},\Theta^{\tau})$ (A.1)
+$=\sum_{y_{1}}\cdots\sum_{y_{N}}\sum_{i=1}^{N}\sum_{k=1}^{K}\delta_{k=y_{i}}\log(\nu_{k}\Pr(\xi_{i}\mid\pi_{r_{k}}))\prod_{i^{\prime}=1}^{N}\Pr(y_{i^{\prime}}\mid\xi_{i^{\prime}},\Theta^{\tau})$ (A.2)
+$=\sum_{k=1}^{K}\sum_{i=1}^{N}\log(\nu_{k}\Pr(\xi_{i}\mid\pi_{r_{k}}))\sum_{y_{1}}\cdots\sum_{y_{N}}\delta_{k=y_{i}}\prod_{i^{\prime}=1}^{N}\Pr(y_{i^{\prime}}\mid\xi_{i^{\prime}},\Theta^{\tau})$ (A.3)
+$=\sum_{k=1}^{K}\sum_{i=1}^{N}\log(\nu_{k}\Pr(\xi_{i}\mid\pi_{r_{k}}))\zeta_{ik}^{\tau}$ (by Equation 7)
+$=\sum_{k=1}^{K}\sum_{i=1}^{N}\zeta_{ik}^{\tau}\log(\nu_{k})+\sum_{k=1}^{K}\sum_{i=1}^{N}\zeta_{ik}^{\tau}\log(\Pr(\xi_{i}\mid\pi_{r_{k}}))\text{\,,}$ (A.4)
 where $\delta$ denotes the Kronecker delta function.
-Equation A.4 indicates that $\nu_{k}$ and ${\mathcal{R}}_{k}$ are not interdependent, we can thus optimize them separately in the M-step of EM, leading to the second update equation in Equation 5 trivially.
+Equation A.4 indicates that $\nu_{k}$ and $r_{k}$ are not interdependent, we can thus optimize them separately in the M-step of EM, leading to the second update equation in Equation 6 trivially.
 According to Gibbs’ inequality, the first term of Equation A.4 is maximized if and only if
@@
 for all $\nu\in\Theta$, proving the first update equation in Equation 6. ∎
-A.2 Proof of Theorem 2
+A.2 Proof of Theorem 4.3
 Proof.
-Similar to the proof for Theorem 1, substitute the parameter set $\Theta=\{\Pi,\Lambda;{\mathcal{R}}_{1},\ldots,{\mathcal{R}}_{K}\}$ into the EM update equation (Equation 4) and unroll:
+Similar to the proof for Theorem 4.2, substitute the parameter set $\Theta=\{\Pi,\Lambda;r_{1},\ldots,r_{K}\}$ into the EM update equation (Equation 5) and unroll:
 $\Psi(\Theta,\Theta^{\tau})$ $\coloneqq\sum_{\mathcal{Y}}\mathcal{L}(\Theta\mid\mathcal{D},\mathcal{Y})\Pr(\mathcal{Y}\mid\mathcal{D},\Theta^{\tau})$ (from Equation 5)
-$=\sum_{{\mathcal{Y}}}\log(\pi_{y_{0}}\Pr(\xi_{0}\mid\pi_{{\mathcal{R}}_{y_{0}}})\prod_{i=1}^{N}\Lambda_{y_{i-1}y_{i}}\Pr(\xi_{i}\mid\pi_{{\mathcal{R}}_{y_{i}}}))\Pr({\mathcal{Y}}\mid{\mathcal{D}},\Theta^{\tau})$ (A.6)
+$=\sum_{\mathcal{Y}}\log(\Pi_{y_{0}}\Pr(\xi_{0}\mid\pi_{r_{y_{0}}})\prod_{i=1}^{N}\Lambda_{y_{i-1}y_{i}}\Pr(\xi_{i}\mid\pi_{r_{y_{i}}}))\Pr(\mathcal{Y}\mid\mathcal{D},\Theta^{\tau})$ (A.6)
 $=\sum_{\mathcal{Y}}\log(\Pi_{y_{0}})\Pr(\mathcal{Y}\mid\mathcal{D},\Theta^{\tau})$
 $\quad+\sum_{\mathcal{Y}}\sum_{i=1}^{N}\log(\Lambda_{y_{i-1}y_{i}})\Pr(\mathcal{Y}\mid\mathcal{D},\Theta^{\tau})$
-$\quad+\sum_{{\mathcal{Y}}}\sum_{i=0}^{N}\log(\Pr(\xi_{i}\mid\pi_{{\mathcal{R}}_{y_{i}}}))\Pr({\mathcal{Y}}\mid{\mathcal{D}},\Theta^{\tau})$ (A.7)
+$\quad+\sum_{\mathcal{Y}}\sum_{i=0}^{N}\log(\Pr(\xi_{i}\mid\pi_{r_{y_{i}}}))\Pr(\mathcal{Y}\mid\mathcal{D},\Theta^{\tau})$ (A.7)
 $=\sum_{y_{1}}\cdots\sum_{y_{N}}\sum_{k=1}^{K}\delta_{k=y_{0}}\log(\Pi_{k})\prod_{i^{\prime}=1}^{N}\Pr(y_{i^{\prime}}\mid\xi_{i^{\prime}},\Theta^{\tau})$
 $\quad+\sum_{y_{1}}\cdots\sum_{y_{N}}\sum_{i=1}^{N}\sum_{k=1}^{K}\sum_{l=1}^{K}\delta_{k=y_{i-1},l=y_{i}}\log(\Lambda_{kl})\prod_{i^{\prime}=1}^{N}\Pr(y_{i^{\prime}}\mid\xi_{i^{\prime}},\Theta^{\tau})$
-$\quad+\sum_{y_{1}}\cdots\sum_{y_{N}}\sum_{i=0}^{N}\sum_{k=1}^{K}\delta_{k=y_{i}}\log(\Pr(\xi_{i}\mid\pi_{{\mathcal{R}}_{k}}))\prod_{i^{\prime}=1}^{N}\Pr(y_{i^{\prime}}\mid\xi_{i^{\prime}},\Theta^{\tau})$ (A.8)
+$\quad+\sum_{y_{1}}\cdots\sum_{y_{N}}\sum_{i=0}^{N}\sum_{k=1}^{K}\delta_{k=y_{i}}\log(\Pr(\xi_{i}\mid\pi_{r_{k}}))\prod_{i^{\prime}=1}^{N}\Pr(y_{i^{\prime}}\mid\xi_{i^{\prime}},\Theta^{\tau})$ (A.8)
 $=\sum_{k=1}^{K}\Pr(y_{0}=k\mid\mathcal{D},\Theta^{\tau})\log(\Pi_{k})$
 $\quad+\sum_{k=1}^{K}\sum_{l=1}^{K}\sum_{i=1}^{N}\Pr(y_{i-1}=k,y_{i}=l\mid\mathcal{D},\Theta^{\tau})\log(\Lambda_{kl})$
-$\quad+\sum_{k=1}^{K}\sum_{i=0}^{N}\Pr(y_{i}=k\mid{\mathcal{D}},\Theta^{\tau})\log(\Pr(\xi_{i}\mid\pi_{{\mathcal{R}}_{k}}))\text{\,,}$ (A.9)
-where $\delta$ denotes the Kronecker delta function.
-Since $\pi_{k}$, $\Lambda_{kl}$ and ${\mathcal{R}}_{k}$ are not interdependent, we can thus maximize the respective term separately, resulting in Equation 6.
-∎
-Remark 1.
-In a more practical case where the agent demonstration space has multiple trajectory sequences, Theorem 2 can also be generalized and proved in the same manner.
-Remark 2.
+$\quad+\sum_{k=1}^{K}\sum_{i=0}^{N}\Pr(y_{i}=k\mid\mathcal{D},\Theta^{\tau})\log(\Pr(\xi_{i}\mid\pi_{r_{k}}))\text{\,,}$ (A.9)
+where $\delta$ denotes the Kronecker delta function. Since $\Pi_{k}$, $\Lambda_{kl}$ and $r_{k}$ are not interdependent, we can thus maximize the respective term separately, resulting in Equation 8. ∎
+Remark A.1.
+In a more practical case where the agent demonstration space has multiple trajectory sequences, Theorem 4.3 can also be generalized and proved in the same manner.
+Remark A.2.
 All $\xi\in\mathcal{D}$ above are assumed to be the trajectory for a whole episode.
-In some special cases where it is assumed that the latent state transition happens after each action execution, instead of per episode, Theorem 2 can also be applied by regarding each episode as a trajectory sequence with each trajectory consists of only one action execution.
+In some special cases where it is assumed that the latent state transition happens after each action execution, instead of per episode, Theorem 4.3 can also be applied by regarding each episode as a trajectory sequence with each trajectory consists of only one action execution.
 Appendix B Algorithms
@@
 The pseudo code for LV-IQL can be found at Algorithm 1.
+Algorithm 1 Latent Variable Inverse Q-learning (LV-IQL)
   Input: agent demonstrations $\mathcal{D}$, latent space dimension $K$
-1 initialize $\Theta\coloneqq\{\nu_{1},\ldots,\nu_{K};{\mathcal{R}}_{1},\ldots,{\mathcal{R}}_{K}\mid\nu_{1}+\cdots+\nu_{K}=1\}$ repeat
-2       E-step
-3             foreach $\xi_{i}\in{\mathcal{D}}$ do
-4                   forall $k$ do
-5                         $\zeta_{ik}\leftarrow\prod_{(s,a)\in\xi_{i}}\pi_{{\mathcal{R}}_{k}}(s,a)\nu_{k}/Z$
-6
-7
-8
-9      M-step
-10             forall $k$ do
-11                   $\nu_{k}\leftarrow\sum_{i=1}^{N}\zeta_{ik}/N$ compute ${\mathcal{R}}_{k}$ via IAVI or IQL on ${\mathcal{D}}$ with weight $\zeta_{ik}$ on trajectory $\xi_{i}$
-12
-13
-14until convergence;
+  Initialize $\Theta\coloneqq\{\nu_{1},\ldots,\nu_{K};r_{1},\ldots,r_{K}\mid\nu_{1}+\cdots+\nu_{K}=1\}$.
+  repeat
+     E-step:
+     for each $\xi_{i}\in\mathcal{D}$ do
+        for all $k$ do
+           $\zeta_{ik}\leftarrow\prod_{(s,a)\in\xi_{i}}\pi_{r_{k}}(s,a)\nu_{k}/Z$
+        end for
+     end for
+     M-step:
+     for all $k$ do
+        $\nu_{k}\leftarrow\sum_{i=1}^{N}\zeta_{ik}/N$
+        Compute $r_{k}$ via IAVI or IQL on $\mathcal{D}$ with weight $\zeta_{ik}$ on trajectory $\xi_{i}$.
+     end for
+  until convergence
   Output: $\Theta$
-Algorithm 1 Latent Variable Inverse Q-learning (LV-IQL)
 B.2 Latent Markov Variable Inverse Q-learning
@@
 Thus the update equation in Equation 8 is equivalent to
-$&\Pi_{k}^{\tau+1}\coloneqq\mathfrak{g}_{0k}^{\tau}\\ &\Lambda_{kl}^{\tau+1}\coloneqq\frac{\sum_{i=1}^{N}\mathfrak{x}_{ikl}^{\tau}}{\sum_{i=0}^{N-1}\mathfrak{g}_{ik}^{\tau}}\\ &{\mathcal{R}}_{k}^{\tau+1}\coloneqq\operatorname*{arg\,max}_{{\mathcal{R}}_{k}}\sum_{i=0}^{N}\mathfrak{g}_{ik}^{\tau}\log(\Pr(\xi_{i}\mid\Pi_{{\mathcal{R}}_{k}}))\text{\,.}\end{aligned}\right.$ (B.5)
+$&\Pi_{k}^{\tau+1}\coloneqq\mathfrak{g}_{0k}^{\tau}\\ &\Lambda_{kl}^{\tau+1}\coloneqq\frac{\sum_{i=1}^{N}\mathfrak{x}_{ikl}^{\tau}}{\sum_{i=0}^{N-1}\mathfrak{g}_{ik}^{\tau}}\\ &r_{k}^{\tau+1}\coloneqq\operatorname*{argmax}_{r_{k}}\sum_{i=0}^{N}\mathfrak{g}_{ik}^{\tau}\log(\Pr(\xi_{i}\mid\pi_{r_{k}}))\text{\,.}\end{aligned}\right.$ (B.5)
 Combining Equation B.5 and the class of IQL algorithms leads to LMV-IQL (Algorithm 2).
-Input: expert demonstrations ${\mathcal{D}}$, latent space dimension $K$
-1 initialize $\Theta\coloneqq\{\Pi,\Lambda;{\mathcal{R}}_{1},\ldots,{\mathcal{R}}_{K}\}$ repeat
-2       E-step
-3             calculate $\mathfrak{g}$ and $\mathfrak{x}$ according to Equations B.1B.4
-4
-5      M-step
-6             forall $k$ do
-7                   $\Pi_{k}\leftarrow\mathfrak{g}_{0k}$ forall $l$ do
-8                         $\Lambda_{kl}\leftarrow\sum_{i=1}^{N}\mathfrak{x}_{ikl}/\sum_{i=0}^{N-1}\mathfrak{g}_{ik}$
-9                  compute ${\mathcal{R}}_{k}$ via IAVI or IQL on ${\mathcal{D}}$ with weight $\mathfrak{g}_{ik}$ on trajectory $\xi_{i}$
-10
-11
-12until convergence;
+Algorithm 2 Latent Markov Variable Inverse Q-learning (LMV-IQL)
+  Input: agent demonstrations $\mathcal{D}$, latent space dimension $K$
+  Initialize $\Theta\coloneqq\{\Pi,\Lambda;r_{1},\ldots,r_{K}\}$.
+  repeat
+     E-step:
+     Calculate $\mathfrak{g}$ and $\mathfrak{x}$ according to Equations B.1 to B.4.
+     M-step:
+     for all $k$ do
+        $\Pi_{k}\leftarrow\mathfrak{g}_{0k}$
+        for all $l$ do
+           $\Lambda_{kl}\leftarrow\sum_{i=1}^{N}\mathfrak{x}_{ikl}/\sum_{i=0}^{N-1}\mathfrak{g}_{ik}$
+        end for
+        Compute $r_{k}$ via IAVI or IQL on $\mathcal{D}$ with weight $\mathfrak{g}_{ik}$ on trajectory $\xi_{i}$.
+     end for
+  until convergence
   Output: $\Theta$
-Algorithm 2 Latent Markov Variable Inverse Q-learning (LMV-IQL)
 Appendix C Further Details and Additional Results on the Simulated Gridworld Behavior Dataset
@@
 ‘Thirsty’ $\bm{0.00\pm 0.00}$ $\bm{2.00\pm 1.79}$ $7.98\pm 0.00$ $7.98\pm 0.00$
-Figure 6: (Top) Visualization of the normalized ground truth and learnt state-value functions.
+Figure C.1: (Top) Visualization of the normalized ground truth and learnt state-value functions.
 Red lines indicate the ground truth trajectory distribution and the learnt trajectory clusters used to recover the reward function for respective intention.
@@
 replacing the $-1$ reward on the type of reward irrelevant to the intentions with 00.
-In this environment, there is an increased overlapping between some of the demonstrated trajectories under different intentions (Figure 6).
+In this environment, there is an increased overlapping between some of the demonstrated trajectories under different intentions (Figure C.1).
 However, LV-IAVI and LV-IQL still outperform the single intention algorithm IAVI and IQL in trajectory clustering and recovering corresponding expert reward functions.
@@
 […]
-Figure 7: (A) Comparison of LMV-IAVI on test set trajectories to a random policy and DIRL, represented as LL.
+Figure D.1: (A) Comparison of LMV-IAVI on test set trajectories to a random policy and DIRL, represented as LL.
 (B) Learnt policy (red arrows) under different intentions and corresponding state occupancy (grey colormap) in the environment.
@@
 In contrast to the outcomes from the water-restricted animal dataset (Figure 3), LMV-IAVI demonstrates a higher test LL, even when considering a single intention.
-As the number of latent states (associated with DIRL’s goal maps) grows, the test LL of LMV-IAVI increases, while the test LL of DIRL remains constant (Figure 7).
+As the number of latent states (associated with DIRL’s goal maps) grows, the test LL of LMV-IAVI increases, while the test LL of DIRL remains constant (Figure D.1).
 Focus on the two latent states LMV-IAVI, the inferred policy under two intentions exhibits ‘Exploring’ and ‘Tired’ behavior.
@@
 That is, the training and test set consisted of 80% and 20% of recorded sessions of each mouse, respectively., and Algorithm 3 was fit on each cross-validation fold independently.
+Algorithm 3 Fitting LMV-IQL on real mice behavior
   Fit IQL globally:
-1 foreach $\ell_{h}\in\{1,\ldots,5\}$ do
-2       run IQL on the concatenated data from all animals until convergence
-select best $\ell_{h}$ with largest test set LL Fit LMV-IQL globally:
-3 foreach $K\in\{2,\ldots,5\}$ do
-4       foreach $i\in\{1,\ldots,10\}$ do
-5             initialize LMV-IQL with $K$ latent states and random parameters run Algorithm 2 on the concatenated data from all animals until convergence
-6
+  for each $\ell_{h}\in\{1,\ldots,5\}$ do
+     Run IQL on the concatenated data from all animals until convergence.
+  end for
+  Select best $\ell_{h}$ with largest test set log-likelihood.
+  Fit LMV-IQL globally:
+  for each $K\in\{2,\ldots,5\}$ do
+     for each $i\in\{1,\ldots,10\}$ do
+        Initialize LMV-IQL with $K$ latent states and random parameters.
+        Run Algorithm 2 on the concatenated data from all animals until convergence.
+     end for
+  end for
   Fit separate LMV-IQL to each animal:
-7 forall animals do
-8       foreach $K\in\{1,\ldots,5\}$ do
-9             initialize LMV-IQL with $K$ latent states using the best global fit parameters for this $K$ run Algorithm 2 until convergence
-10
-Algorithm 3 Fitting LMV-IQL on real mice behavior
-fBwB4ngcAcByHJpPJl+fn54mD3Gg0NrquXxeLRQAAwzAYj8cwTZPwPH9/sVg8PXweDAauqqr2cDjEer1GJBLBZDJBs9mE4zjwfZ85lAGg2+06hmGgXq+j3+/DsixYlgVN03a9Xu8jgCNCyIegIAgx13Vfd7vdu+FweG8YRkjXdWy329+dTgeSJD3ieZ7RNO0VAXAPwDEAO5VKndi2fWrb9jWl9Esul6PZbDY9Go1OZ7PZ9z/lyuD3OozU2wAAAABJRU5ErkJggg==" alt="[LOGO]">
+  for all animals do
+     for each $K\in\{1,\ldots,5\}$ do
+        Initialize LMV-IQL with $K$ latent states using the best global fit parameters for this $K$.
+        Run Algorithm 2 until convergence.
+     end for
+  end for